### v2

# Existence and Uniqueness for a Porous-Medium-Type Diffusion Model with Time-Dependent Growth

Abstract
This paper investigates a nonlinear diffusion equation, characterized by a power-law dispersal mechanism and a time-dependent growth rate in a bounded domain with homogeneous Dirichlet boundary conditions. We establish three main theoretical contributions: first, the existence and uniqueness of separable classical solutions via a transformation into a stationary sublinear elliptic framework; second, a generalized comparison principle with respect to both initial conditions and growth rates; and third, a general existence and uniqueness result for non-separable solutions via a monotone iteration method. The connection between the temporal dynamics and the spatial profile is made explicit, providing analytical expressions for the solution. Furthermore, we demonstrate the model’s effectiveness as a successful application in digital image denoising, showing superior robustness compared to the classical Perona–Malik model.
Keywords: Nonlinear diffusion; Biological populations; Gurtin–MacCamy model; Porous medium equation; Existence and uniqueness; Monotone iteration; Numerical simulation; Bernoulli equation.
MSC 2020: 35K55; 35K57; 35J20; 35B40; 35D30; 92D25.
1. Introduction
In this work, we focus on the nonlinear diffusion equation:
$\frac{\partial v}{\partial t}(x,t)-\Delta v^{\alpha}(x,t)=\mu\left(t\right)v(x,t),\quad\text{in }\Omega\times(0,T),$ (1.1)
where $\alpha>1$, $T>0$ (possibly infinite), $\mu:\left[0,\infty\right)\rightarrow\left(0,\infty\right)$ is a continuous time-dependent growth rate, and $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded smooth domain, with hostile boundaries ($\partial\Omega$).
This equation, denoted as (1.1) models the spatial spread of biological populations. Here, $v(x,t)$ represents the evolving population density (or, analogously, temperature or concentration), and the term $\Delta v^{\alpha}(x,t)$ embodies a power-law diffusion process that underpins both biological dispersion and energy transport in complex materials. The parameter $\alpha$ governs the nonlinearity level in the diffusion process, $\mu(t)$ represents the population growth rate which depends on time, and $T$ defines the observation time frame.
Problem (1.1) is supplemented with boundary and initial conditions:
$v(x,t)$ $=0,\qquad\text{on }\partial\Omega\times(0,T),$ (1.2)
$v(x,0)$ $=v_{0}(x),\qquad\text{on }\Omega.$ (1.3)
It is well known that problem (1.1), along with boundary conditions (1.2)-(1.3), is frequently studied within the framework established by Gurtin and MacCamy [8]. The boundary condition (1.2) ensures the solution vanishes at the boundary of the domain, while the initial condition (1.3) specifies the initial state of the population.
Additionally, studying the coupled problem (1.1) -(1.2)-(1.3) is connected to the asymptotic behavior of the solution to
$\left\{\begin{array}[]{ccc}-\Delta u\left(x\right)=u^{\frac{1}{\alpha}}\left(x\right)&in&\Omega,\\ u\left(x\right)>0&in&\Omega,\\ u\left(x\right)=0&on&\partial\Omega.\end{array}\right$. (1.4)
A key methodological feature of our study is the separation-of-variables technique, which allows us to express solutions of (1.1) -(1.2)-(1.3) in the form
$v_{\gamma}\left(x,t\right)=S\left(t\right)u^{\frac{1}{\alpha}}\left(x\right),\quad S\left(t\right)=\left[e^{\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\left(\gamma^{1-\alpha}+(\alpha-1)\int_{0}^{t}e^{-\left(1-\alpha\right)\int_{0}^{\tau}\mu\left(s\right)ds}d\tau\right)\right]^{\frac{1}{1-\alpha}},$ (1.5)
thereby decoupling temporal dynamics from spatial behavior. Here, $u\left(x\right)$ is a solution of (1.4) and the initial condition is satisfied as
$v_{0}(x)=v_{\gamma}\left(x,0\right)=S\left(0\right)u^{\frac{1}{\alpha}}\left(x\right),$ (1.6)
where $S\left(0\right)=\gamma\in\left(0,\infty\right)$ is given.
In biological systems, nonlinear diffusion equations are employed to capture the complex dynamics of microbial populations, where growth and dispersal interact in nontrivial ways [8, 10, 9]. At the same time, similar mathematical models are vital in describing thermal processes in non-homogeneous materials, as well as various phenomena in material science such as phase transitions and the evolution of microstructural defects [14, 6].
The mathematical theory of nonlinear diffusion equations, particularly those of porous medium type, has a rich history. The foundational work of Gurtin and MacCamy [8] introduced the density-dependent diffusion model for biological populations, where the diffusion coefficient depends on the population density itself. This leads to equations of the form $v_{t}=\Delta v^{\alpha}$ with $\alpha>1$, known as the slow diffusion case. The monograph by Vázquez [14] provides a comprehensive treatment of the mathematical theory for such equations, including existence, uniqueness, regularity, and asymptotic behavior. The work of Aronson and Peletier [1] established fundamental results on the large-time behavior of solutions in bounded domains.
The elliptic counterpart of our problem, namely the sublinear equation $\Delta u=u^{1/\alpha}$ with $\alpha>1$, has been extensively studied. The celebrated work of Brezis and Oswald [3] established existence and uniqueness results for sublinear elliptic problems, while Brezis and Kamin [2] extended these results to unbounded domains. These elliptic results are fundamental to our approach, as they provide the spatial profile for the separable solutions.
Recent advances in the analysis and numerical treatment of diffusion-type equations highlight the vibrant research landscape in this field. The subsolution–supersolution method, systematically developed by Pao [11] for parabolic systems, provides a powerful framework for establishing existence and uniqueness in the general non-separable case. Parallel to these technical developments, recent progress on population models with nonlinear or nonstandard survival structures has been reported; see, for example, the pseudo-exponential survival framework in [5] and the analysis of equilibrium solutions in a nonlinear separable setting in [4]. These contributions collectively underscore the importance of developing robust theoretical and computational frameworks for complex diffusion phenomena, setting the stage for our investigation into the nonlinear model with time-dependent growth.
The broader applicability of this framework extends well beyond biological dynamics. In thermal diffusion, such nonlinear models describe the distribution of temperature in materials where conduction properties vary with the temperature itself, leading to power-law behaviors. Similarly, in material sciences, these equations underpin models of phase transitions and the migration of defects, offering insights into microstructural stability.
In summary, this work establishes two key theoretical results complemented by rigorous numerical verification:
  • Theorem 1.1: An existence and uniqueness result for separable solutions, obtained by transforming the parabolic problem into an elliptic equation coupled with a Bernoulli-type ordinary differential equation.
  • Theorem 1.2: A comparison principle result with respect to the initial conditions, showing that the population distribution preserves its initial ordering throughout the evolution.
  • Theorem 1.3: A general existence and uniqueness result for non-separable solutions, established via the monotone iteration method based on subsolution–supersolution techniques.
These results are presented below in an abstract formulation. The first result is the following:
Theorem 1.1.
Let $\Omega\subset\mathbb{R}^{N}$ be a bounded, arcwise-connected open set with a compact closure and boundary $\partial\Omega$ of class $C^{3}$. Under these conditions, the nonlinear diffusion problem (1.1)–(1.2) admits a unique solution $v_{\gamma}(x,t)$, given by (1.5), where $u\in C^{2}(\Omega)\cap C^{1}(\overline{\Omega})$ exists uniquely, and the initial boundary condition (1.3) is given by (1.6).
Our methodology for proving Theorem 1.1 departs from classical strategies, such as those found in [2, 8, 15], by harnessing a transformation that simplifies the analysis even under challenging regimes. This approach not only strengthens the theoretical framework but also paves the way for robust numerical implementations. Next, we proceed with a comparison principle.
Theorem 1.2.
Let $v_{0},w_{0}\in C(\overline{\Omega})$ be two initial conditions such that $v_{0}(x)\leq w_{0}(x)$ for all $x\in\overline{\Omega}$. Assume that $\mu_{1}\left(t\right)$ and $\mu_{2}\left(t\right)$ are continuous and bounded functions such that $\mu_{1}\left(t\right)\leq\mu_{2}\left(t\right)$. If $v(x,t)$ and $w(x,t)$ are the unique classical solutions to problem (1.1)–(1.3) corresponding to $(v_{0},\mu_{1})$ and $(w_{0},\mu_{2})$, respectively, then
$v(x,t)\leq w(x,t)\quad\text{for all }\left(x,t\right)\in\overline{\Omega}\times[0,T).$ (1.7)
Intuitively, this result shows that if the initial population $v_{0}$ is smaller than $w_{0}$ and the growth rate $\mu_{1}(t)$ is smaller than $\mu_{2}(t)$, then the solution $v$ stays below $w$ during the entire observation period.
Theorem 1.1 and Theorem 1.2 leads to a new result which is presented here for the first time.
Theorem 1.3.
Let $\Omega\subset\mathbb{R}^{N}$ be a bounded, arcwise-connected open set with a compact closure and boundary $\partial\Omega$ of class $C^{3}$. Assume that $u_{-},u_{+}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ satisfy
$-\Delta u_{-}$ $\leq u_{-}^{1/\alpha}\qquad\text{in }\Omega,$ $u_{-}>0\text{ in }\Omega,$ $u_{-}=0\text{ on }\partial\Omega,$ (1.8)
$-\Delta u_{+}$ $\geq u_{+}^{1/\alpha}\qquad\text{in }\Omega,$ $u_{+}>0\text{ in }\Omega,$ $u_{+}=0\text{ on }\partial\Omega,$ (1.9)
and $u_{-}\leq u_{+}$ in $\overline{\Omega}$. Fix $\gamma>0$ and let $S$ be the unique solution of
$S^{\prime}(t)+S^{\alpha}(t)\,=\mu\left(t\right)S\left(t\right),\qquad S(0)=\gamma,$ (1.10)
given by (1.5). Define $v_{-}(x,t)=S(t)u_{-}^{1/\alpha}(x)$ and $v_{+}(x,t)=S(t)u_{+}^{1/\alpha}(x)$. If $v_{0}\in C(\overline{\Omega})$ satisfies $v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)$ in $\overline{\Omega}$, then there exists a unique classical solution $v(x,t)$ to (1.1)-(1.3), such that $v_{-}(x,t)\leq v(x,t)\leq v_{+}(x,t)$ for all $\left(x,t\right)\in\overline{\Omega}\times\left[0,T\right)$.
A careful examination of Theorems 1.1 and 1.3 and their proofs shows that both results play an essential role, even though Theorem 1.1 is a particular case of Theorem 1.3. Theorem 1.1 remains significant because its proof lends itself to straightforward implementation in a programming language through simplified numerical procedures, whereas Theorem 1.3 applies to the more general situation in which the solution is not necessarily separable. As an illustration, we present a well-known result that highlights the distinct roles played by Theorems 1.1 and 1.3.
Remark 1.4.
There always exists a continuous function
$u_{-}^{+}:\overline{\Omega}\rightarrow[0,\infty)$
such that
$u_{-}(x)\leq u_{-}^{+}(x)\leq u_{+}(x)\quad\text{for all }x\in\overline{\Omega}.$ (1.11)
For instance, for any continuous function $\theta:\overline{\Omega}\rightarrow[0,1]$, the function
$u_{-}^{+}(x):=(1-\theta(x))\,u_{-}(x)+\theta(x)\,u_{+}(x)$
satisfies (1.11).
Therefore, Theorem 1.2 strengthens both the first (Theorem 1.1) and the third result (Theorem 1.3). First, it plays a key role in the proof of Theorem 1.3; second, when combined with Theorem 1.1, it enables the construction of a subsolution–supersolution pair, which in turn makes the implementation of the results in a computer language feasible.
The remainder of the paper is structured as follows. Section 2 provides the proofs of Theorems 1.11.3. In Section 3, we derive the model from conservation laws and energy considerations, following the framework of Gurtin and MacCamy. Section 4 presents conclusions and discusses future research directions. Finally, Appendices AD provide the complete Python implementations and present the studied model as a successful application in digital image processing, for clarity and reproducibility.
2. Proof of the Main Results
In the following proof of the Theorem 1.1, we apply the successive approximation method, as our goal is to implement the result using a computer language.
Proof of Theorem 1.1.
We first prove the existence of solution to the problem (1.4). Let $\phi_{1}\in C^{2}(\overline{\Omega})$ be the first positive eigenfunction corresponding to the first eigenvalue $\lambda_{1}$ of the problem
$\left\{\begin{array}[]{lll}-\Delta\phi(x)=\lambda\phi(x),&in&\Omega,\\ \phi(x)=0,&on&\partial\Omega,\end{array}\right$.
and $w\in C^{2}(\overline{\Omega})$ be the unique solution of the problem
$\left\{\begin{array}[]{ccc}-\Delta w(x)=1,&in&\Omega,\\ w(x)=0,&on&\partial\Omega.\end{array}\right$.
We choose
$\sigma=\frac{1}{\lambda_{1}^{\frac{\alpha}{\alpha-1}}\max_{x\in\overline{\Omega}}\phi_{1}\left(x\right)}\text{ and }c=\max\left\{\sigma\lambda_{1}\max_{x\in\overline{\Omega}}\phi_{1},\left(\max_{x\in\overline{\Omega}}w\left(x\right)\right)^{\frac{1}{\alpha-1}}\right\}.$
Then, the function $\underline{u}(x)=\sigma\phi_{1}$ is a subsolution of (1.4) while $\overline{u}\left(x\right)=cw(x)$ is a supersolution. Indeed,
$-\Delta\underline{u}(x)-\left(\underline{u}(x)\right)^{\frac{1}{\alpha}}$ $=\sigma\lambda_{1}\phi_{1}\left(x\right)-\left(\sigma\phi_{1}\right)^{\frac{1}{\alpha}}$
$=\sigma\lambda_{1}\phi_{1}\left(x\right)-\sigma^{\frac{1}{\alpha}}\left(\phi_{1}\left(x\right)\right)^{\frac{1}{\alpha}}$
$=\sigma\left(\lambda_{1}\phi_{1}\left(x\right)\right)^{\frac{1}{\alpha}}\left[\left(\lambda_{1}\phi_{1}\left(x\right)\right)^{\frac{\alpha-1}{\alpha}}-\frac{1}{\sigma}\left(\frac{\sigma}{\lambda_{1}}\right)^{\frac{1}{\alpha}}\right]$
$\leq\sigma\left(\lambda_{1}\phi_{1}\left(x\right)\right)^{\frac{1}{\alpha}}\left[\left(\lambda_{1}\max_{x\in\overline{\Omega}}\phi_{1}\left(x\right)\right)^{\frac{\alpha-1}{\alpha}}-\frac{1}{\sigma}\left(\frac{\sigma}{\lambda_{1}}\right)^{\frac{1}{\alpha}}\right]=0$
and
$-\Delta\overline{u}(x)-\left(\overline{u}\left(x\right)\right)^{\frac{1}{\alpha}}$ $=c-\left(cw\left(x\right)\right)^{\frac{1}{\alpha}}\geq c-\left(c\max_{x\in\overline{\Omega}}w\left(x\right)\right)^{\frac{1}{\alpha}}$
$=c^{\frac{1}{\alpha}}\left[c^{\frac{\alpha-1}{\alpha}}-\left(\max_{x\in\overline{\Omega}}w\left(x\right)\right)^{\frac{1}{\alpha}}\right]\geq 0.$
Now, since
$\left\{\begin{array}[]{lll}-\Delta[\overline{u}(x)-\underline{u}(x)]=c-\sigma\lambda_{1}\phi_{1}\geq c-\sigma\lambda_{1}\max_{x\in\overline{\Omega}}\phi_{1}\geq 0&in&\Omega\\ \overline{u}(x)-\underline{u}(x)=0&on&\partial\Omega\end{array}\right$.
it follows from the maximum principle that $\underline{u}(x)\leq\overline{u}(x)$, for all $x\in\overline{\Omega}$.
Next, we define the sequence $\{u^{k}\}_{k\geq 0}$ by
$u^{0}(x)=u_{0}(x),\text{ }-\Delta u^{k}\left(x\right)=\left(u^{k-1}\left(x\right)\right)^{\frac{1}{\alpha}}\text{ for }k=1,2,...\text{ and }x\in\Omega$ (2.1)
where
$u_{0}(x)=\underline{u}(x)\quad\forall\,x\in\overline{\Omega},$
with $\underline{u}(x)$ being a subsolution of the problem
$\Delta u(x)=u(x)^{\frac{1}{\alpha}},\quad u(x)>0\text{ in }\Omega,\qquad u(x)=0\text{ on }\partial\Omega.$ (2.2)
Using the Brezis-Oswald inequality [3, Theorems 1 and 2], we derive
$\underline{u}(x)\leq u^{0}(x)\leq u^{1}(x)\leq u^{2}(x)\leq\cdots\leq\overline{u}(x),\quad\forall\,x\in\overline{\Omega}.$ (2.3)
Since the sequence $\{u^{k}\}_{k\geq 0}$ is monotone increasing and bounded above (by $\overline{u}(x)$), it follows by the monotone convergence theorem that the pointwise limit
$u(x)=\lim_{k\rightarrow\infty}u^{k}(x),$
exists for all $x\in\overline{\Omega}$. Standard regularity theory, which can be found in [7], ensures that $u(x)\in C^{2}(\Omega)\cap C^{1}(\overline{\Omega})$ and, by construction, satisfies the required conditions for problem (1.4).
$\underline{u}(x)\leq u(x)\leq\overline{u}(x)\quad\forall\,x\in\overline{\Omega}.$
The uniqueness of the solution to the problem (1.4) follow again from the results established by Brezis and Oswald [3, Theorems 1 and 2] (see also [1] or [2]).
Now, let us prove the validity of (1.5). Assume a solution of the form
$v(x,t)=u^{\frac{1}{\alpha}}(x)\,S(t).$ (2.4)
Then we have
$v^{\alpha}(x,t)=u(x)\,S^{\alpha}(t).$
Substitute (2.4) into (1.1) to obtain:
$u^{\frac{1}{\alpha}}(x)\,S^{\prime}(t)-S^{\alpha}(t)\,\Delta u\left(x\right)=\mu\left(t\right)u^{\frac{1}{\alpha}}(x)\,S(t).$ (2.5)
Dividing (2.5) by $u^{\frac{1}{\alpha}}(x)\,$, we have
$S^{\prime}(t)-S^{\alpha}(t)\,\frac{\Delta u\left(x\right)}{u^{\frac{1}{\alpha}}(x)}=\mu\left(t\right)S\left(t\right),$ (2.6)
or, equivalently
$S^{\prime}(t)+S^{\alpha}(t)\,=\mu\left(t\right)S\left(t\right),$ (2.7)
a Bernoulli-type equation. Multiplying both sides of (2.7) by $S^{-\alpha}(t)$ yields
$S^{\prime}(t)S^{-\alpha}(t)\,=\mu\left(t\right)S^{1-\alpha}\left(t\right)-1.$
We now solve this equation by the standard substitution. Define
$z(t)=S(t)^{\,1-\alpha}.$ (2.8)
Then, by differentiating in (2.8), we have
$z^{\prime}(t)=\left(1-\alpha\right)S^{-\alpha}(t)S^{\prime}\left(t\right)\Longrightarrow S^{-\alpha}(t)S^{\prime}\left(t\right)=\frac{z^{\prime}(t)}{1-\alpha}.$
Substitute the expression for $S^{-\alpha}(t)S^{\prime}\left(t\right)$ into the above:
$\frac{z^{\prime}(t)}{1-\alpha}=\mu\left(t\right)z(t)-1.$
This simplifies to
$z^{\prime}(t)=\left(1-\alpha\right)\mu\left(t\right)z(t)+\alpha-1.$ (2.9)
Multiplying both sides of (2.9) by $e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}$ yields
$\left(z\left(t\right)e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\right)^{\prime}=\left(\alpha-1\right)e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds},$
from where
$z\left(t\right)=e^{\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\int\left(\alpha-1\right)e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}dt.$
But recall that $S(t)^{\,1-\alpha}=z(t)$; hence,
$S\left(t\right)=\left[e^{\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\left(\gamma^{1-\alpha}+(\alpha-1)\int_{0}^{t}e^{-\left(1-\alpha\right)\int_{0}^{\tau}\mu\left(s\right)ds}d\tau\right)\right]^{\frac{1}{1-\alpha}}.$ (2.10)
Replacing (2.10) in (2.4) we obtain the form (1.5). This concludes the proof of Theorem 1.1. ∎
We proceed by establishing the proof of the comparison principle.
Proof of Theorem 1.2.
Let $\mathcal{L}_{\mu}[u]:=\frac{\partial u}{\partial t}-\Delta u^{\alpha}-\mu(t)u$ be the parabolic operator associated with (1.1). By hypothesis, $v$ and $w$ are the unique classical solutions satisfying $\mathcal{L}_{\mu_{1}}[v]=0$ and $\mathcal{L}_{\mu_{2}}[w]=0$ in $\Omega\times(0,T)$. We first establish that $v$ serves as a subsolution for the operator $\mathcal{L}_{\mu_{2}}$. Indeed, substituting $v$ into $\mathcal{L}_{\mu_{2}}$ yields:
$\mathcal{L}_{\mu_{2}}[v]=\frac{\partial v}{\partial t}-\Delta v^{\alpha}-\mu_{2}(t)v=\mu_{1}(t)v-\mu_{2}(t)v=(\mu_{1}(t)-\mu_{2}(t))v.$ (2.11)
Since $\mu_{1}(t)\leq\mu_{2}(t)$ for all $t$ and $v(x,t)\geq 0$ in $\Omega\times(0,T)$, it follows that $(\mu_{1}(t)-\mu_{2}(t))v\leq 0$, which implies:
$\frac{\partial v}{\partial t}-\Delta v^{\alpha}\leq\mu_{2}(t)v\quad\text{in }\Omega\times(0,T).$ (2.12)
On the other hand, the solution $w$ satisfies:
$\frac{\partial w}{\partial t}-\Delta w^{\alpha}=\mu_{2}(t)w\quad\text{in }\Omega\times(0,T).$ (2.13)
Furthermore, the initial and boundary conditions satisfy $v(x,0)=v_{0}(x)\leq w_{0}(x)=w(x,0)$ in $\Omega$ and $v(x,t)=0=w(x,t)$ on $\partial\Omega\times(0,T)$.
To establish the ordering $v\leq w$, we define z(x,t):=v(x,t)w(x,t)z(x,t):=v(x,t)-w(x,t) and note that $z$ satisfies:
$\frac{\partial z}{\partial t}-\Delta(v^{\alpha}-w^{\alpha})=\mu_{1}(t)v-\mu_{2}(t)w\leq\mu_{2}(t)(v-w)=\mu_{2}(t)z.$ (2.14)
Using the monotonicity of the mapping $s\mapsto s^{\alpha}$ for $\alpha>1$ and $s\geq 0$, we can express the difference $v^{\alpha}-w^{\alpha}$ as c(x,t)(vw)c(x,t)(v-w) where $c(x,t)=\int_{0}^{1}\alpha[w+\theta(v-w)]^{\alpha-1}d\theta\geq 0$. Substituting this into the inequality for $z$, we obtain:
$\frac{\partial z}{\partial t}-\Delta(c(x,t)z)-\mu_{2}(t)z\leq 0.$ (2.15)
Given that $z(x,0)\leq 0$ and $z$ vanishes on the boundary, the application of the maximum principle for porous-medium-type equations (see [14, Chapter 3] or [11, Theorem 2.1]) ensures that $z(x,t)\leq 0$ on $\overline{\Omega}\times[0,T)$. Consequently, $v(x,t)\leq w(x,t)$ for all $(x,t)\in\overline{\Omega}\times[0,T)$, which completes the proof. ∎
We end the section by establishing the proof of the third main result.
Proof of Theorem 1.3.
We first compute the quantities appearing in the parabolic operator for the functions $v_{\pm}$.
Step 1: Structure of $v_{\gamma}$. Let $u$ be any of $u_{-}$ or $u_{+}$, and define
$v_{\gamma}(x,t):=S(t)\,u^{1/\alpha}(x).$
Then
$v_{\gamma}^{\alpha}(x,t)=\bigl(S(t)\,u^{1/\alpha}(x)\bigr)^{\alpha}=S(t)^{\alpha}\,u(x).$
Hence, using that $S$ depends only on $t$,
$\Delta\bigl(v_{\gamma}^{\alpha}(x,t)\bigr)=\Delta\bigl(S(t)^{\alpha}u(x)\bigr)=S(t)^{\alpha}\,\Delta u(x).$
Moreover,
$\frac{\partial v_{\gamma}}{\partial t}(x,t)=S^{\prime}\left(t\right)u^{1/\alpha}(x).$
Therefore, the left-hand side of the parabolic operator applied to $v_{\gamma}$ is
$\frac{\partial v_{\gamma}}{\partial t}(x,t)-\Delta\bigl(v_{\gamma}^{\alpha}(x,t)\bigr)-\mu(t)\,v_{\gamma}(x,t)$ $=S^{\prime}\left(t\right)u^{1/\alpha}(x)-S(t)^{\alpha}\,\Delta u(x)-\mu(t)\,S(t)\,u^{1/\alpha}(x)$
$=\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u^{1/\alpha}(x)-S(t)^{\alpha}\,\Delta u(x).$
Step 2: Use of the elliptic inequalities.
(a) Subsolution case. Assume $u=u_{-}$ satisfies (1.8), i.e.
$\Delta u_{-}\leq u_{-}^{1/\alpha}\quad\text{in }\Omega.$
This is equivalent to
$\Delta u_{-}\geq-u_{-}^{1/\alpha}\quad\text{in }\Omega.$
Then, for $v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x)$, we have
$\frac{\partial v_{-}}{\partial t}-\Delta\bigl(v_{-}^{\alpha}\bigr)-\mu(t)\,v_{-}$ $=\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{-}^{1/\alpha}-S(t)^{\alpha}\,\Delta u_{-}$
$\leq\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{-}^{1/\alpha}+S(t)^{\alpha}\,u_{-}^{1/\alpha}$
$=\bigl(-S^{\alpha}(t)\bigr)\,u_{-}^{1/\alpha}+S(t)^{\alpha}\,u_{-}^{1/\alpha}=0,$
where we have used (1.10). Therefore,
$\frac{\partial v_{-}}{\partial t}-\Delta\bigl(v_{-}^{\alpha}\bigr)-\mu(t)\,v_{-}\leq 0\quad\text{in }\Omega\times(0,T),$
which shows that $v_{-}$ is a subsolution of (1.1). On the boundary, since $u_{-}=0$ on $\partial\Omega$, we have
$v_{-}(x,t)=S(t)\,0^{1/\alpha}=0$
on $\partial\Omega\times(0,T)$.
(b) Supersolution case. Assume $u=u_{+}$ satisfies (1.9), i.e.
$\Delta u_{+}\geq u_{+}^{1/\alpha}\quad\text{in }\Omega,$
equivalently
$\Delta u_{+}\leq-u_{+}^{1/\alpha}\quad\text{in }\Omega.$
Then, for $v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x)$, we obtain
$\frac{\partial v_{+}}{\partial t}-\Delta\bigl(v_{+}^{\alpha}\bigr)-\mu(t)\,v_{+}$ $=\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{+}^{1/\alpha}-S(t)^{\alpha}\,\Delta u_{+}$
$\geq\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{+}^{1/\alpha}+S(t)^{\alpha}\,u_{+}^{1/\alpha}$
$=\bigl(-S^{\alpha}(t)\bigr)\,u_{+}^{1/\alpha}+S(t)^{\alpha}\,u_{+}^{1/\alpha}.$
Again, by (1.10), the factor in parentheses vanishes, so
$\frac{\partial v_{+}}{\partial t}-\Delta\bigl(v_{+}^{\alpha}\bigr)-\mu(t)\,v_{+}\geq 0\quad\text{in }\Omega\times(0,T),$
and $v_{+}$ is a supersolution of (1.1). The boundary condition $v_{+}=0$ on $\partial\Omega\times(0,T)$ follows from $u_{+}=0$ on $\partial\Omega$.
Step 3: Ordering of the solution via comparison.
Assume now that $v$ is a classical solution of (1.1) with
$v(x,0)=v_{0}(x)\quad\text{in }\Omega,\qquad v(x,t)=0\quad\text{on }\partial\Omega\times(0,T),$
and that
$v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)\quad\text{for all }x\in\Omega.$
By construction,
$v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x),\qquad v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x),\quad(x,t)\in\overline{\Omega}\times[0,T),$
so the initial data is sandwiched between the sub- and supersolution at the time $t=0$. Moreover, for all $x\in\overline{\Omega}$ yields and $t\in[0,T)$ we have
$v_{-}(x,t)\leq v_{+}(x,t).$
To complete the argument, we now invoke the classical subsolution–supersolution theory for nonlinear parabolic equations. We proceed in several steps.
Step 3.1: Initial ordering. By hypothesis, the initial data $v_{0}$ satisfies
$v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)\quad\text{for all }x\in\overline{\Omega}.$
Since $v_{-}(x,0)=\gamma\,u_{-}^{1/\alpha}(x)$ and $v_{+}(x,0)=\gamma\,u_{+}^{1/\alpha}(x)$, and because $u_{-}\leq u_{+}$ in $\overline{\Omega}$, this ordering is consistent with the assumptions.
Step 3.2: Boundary compatibility. On the boundary $\partial\Omega\times(0,T)$, both $v_{-}$ and $v_{+}$ vanish:
$v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x)=0,\qquad v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x)=0,$
since $u_{-}=u_{+}=0$ on $\partial\Omega$. This ensures that both the subsolution and supersolution are compatible with the homogeneous Dirichlet boundary condition (1.2).
Step 3.3: Application of comparison principles. The theory developed by Pao [11] (see also [8] for related results) establishes that, under suitable regularity assumptions, if $v_{-}$ and $v_{+}$ are respectively an ordered subsolution and supersolution of the parabolic problem (1.1) with boundary conditions (1.2), and if they satisfy the initial ordering at $t=0$, then there exists a unique classical solution $v\in C^{2,1}(\Omega\times(0,T))\cap C(\overline{\Omega}\times[0,T))$ of the problem (1.1)–(1.2)–(1.3) such that
$v_{-}(x,t)\leq v(x,t)\leq v_{+}(x,t)\quad\text{for all }(x,t)\in\overline{\Omega}\times[0,T).$ (2.16)
Step 3.4: Verification of regularity hypotheses. The regularity assumptions required by the theory are satisfied:
  • (i)
    The domain $\Omega$ is of class $C^{3}$ with compact closure;
  • (ii)
    The functions $u_{-},u_{+}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ are strictly positive in $\Omega$ and vanish on $\partial\Omega$;
  • (iii)
    The function $S\in C^{1}([0,T])$ is strictly positive (since $\gamma>0$ and the ODE (1.10) preserves positivity);
  • (iv)
    The initial data $v_{0}\in C(\overline{\Omega})$ is sandwiched between $v_{-}(\cdot,0)$ and $v_{+}(\cdot,0)$.
These conditions ensure that the nonlinear parabolic operator
$\mathcal{L}[v]:=\frac{\partial v}{\partial t}-\Delta(v^{\alpha})-\mu(t)v$
is well-defined and that the subsolution–supersolution method applies.
Step 3.4.1: Monotone iterative construction. To make the existence proof constructive and suitable for numerical implementation, we describe the monotone iteration scheme following Pao [11]. This provides an algorithm for approximating the solution $v$ via sequences that converge monotonically from below and above.
(a) Choice of monotonicity constant. We first choose a function $c(x,t)\geq 0$ sufficiently large to ensure monotonicity. Specifically, consider the nonlinearity in (1.1):
$f(v,t):=\mu(t)v.$
Since $\frac{\partial f}{\partial v}=\mu(t)$, we can take
$c(t):=\max\{0,\,\sup_{s\in[0,t]}|\mu(s)|\},$ (2.17)
where clearly in the case $\mu(s)>0$ we obtain $c(t)=\sup_{s\in[0,t]}\mu(s)$. This ensures that the auxiliary function
$F(v,t):=c(t)\,v+f(v,t)=\bigl(c(t)+\mu(t)\bigr)v$
is nondecreasing in $v$ for $v\in[v_{-},v_{+}]$.
(b) Iterative scheme. Define two sequences $\{v^{(m)}_{-}\}_{m=0}^{\infty}$ and $\{v^{(m)}_{+}\}_{m=0}^{\infty}$ as follows. Initialize with
$v^{(0)}_{-}(x,t):=v_{-}(x,t),\qquad v^{(0)}_{+}(x,t):=v_{+}(x,t).$ (2.18)
For $m\geq 1$, construct $v^{(m)}_{-}$ and $v^{(m)}_{+}$ by solving the linearized parabolic problem
$\frac{\partial v^{(m)}}{\partial t}-\Delta\bigl((v^{(m)})^{\alpha}\bigr)+c(t)\,v^{(m)}$ $=c(t)\,v^{(m-1)}+\mu(t)\,v^{(m-1)}\qquad\text{in }\Omega\times(0,T),$ (2.19)
$v^{(m)}(x,t)$ $=0\qquad\text{on }\partial\Omega\times(0,T),$
$v^{(m)}(x,0)$ $=v_{0}(x)\qquad\text{in }\Omega.$
Here, the superscript $(m)$ denotes the iteration index, and $v^{(m-1)}$ is the previous iterate.
(c) Monotonicity and convergence. By the comparison principle for parabolic equations and the choice of $c(t)$, the sequences satisfy
$v_{-}=v^{(0)}_{-}\leq v^{(1)}_{-}\leq v^{(2)}_{-}\leq\cdots\leq v^{(m)}_{-}\leq\cdots\leq v^{(m)}_{+}\leq\cdots\leq v^{(2)}_{+}\leq v^{(1)}_{+}\leq v^{(0)}_{+}=v_{+}.$ (2.20)
That is, the lower sequence $\{v^{(m)}_{-}\}$ is monotone increasing and bounded above by $v_{+}$, while the upper sequence $\{v^{(m)}_{+}\}$ is monotone decreasing and bounded below by $v_{-}$. Standard parabolic regularity estimates (see [11], Theorem 2.1) ensure that both sequences converge uniformly on compact subsets of $\overline{\Omega}\times[0,T)$ to limits
$\underline{v}(x,t):=\lim_{m\to\infty}v^{(m)}_{-}(x,t),\qquad\overline{v}(x,t):=\lim_{m\to\infty}v^{(m)}_{+}(x,t).$
Moreover, both $\underline{v}$ and $\overline{v}$ are classical solutions of (1.1)–(1.2)–(1.3), and they are the minimal and maximal solutions, respectively, in the ordered interval $v_{-},v_{+}]$.
(d) Uniqueness implies convergence to the same limit. If the solution is unique (as established in Step 3.5), then $\underline{v}\equiv\overline{v}\equiv v$, and both sequences converge to the same solution:
$\lim_{m\to\infty}v^{(m)}_{\pm}(x,t)=v(x,t)\quad\text{uniformly on }\overline{\Omega}\times[0,T).$ (2.21)
This constructive approach is particularly well-suited for numerical implementation: at each iteration $m$, one solves a linear parabolic problem (since the right-hand side depends only on the previous iterate $v^{(m-1)}$), which can be discretized using standard finite difference or finite element methods. The monotone convergence (2.20) provides a natural stopping criterion based on the $L^{\infty}$ or $L^{2}$ norm of the difference $\|v^{(m)}-v^{(m-1)}\|$.
Step 3.5: Uniqueness. Uniqueness of the classical solution follows from the monotonicity properties of the operator. Specifically, if $v^{(1)}$ and $v^{(2)}$ are two solutions both lying in the interval $v_{-},v_{+}]$, then the difference $w:=v^{(1)}-v^{(2)}$ satisfies
$\frac{\partial w}{\partial t}-\text{div}[\nabla(v^{(1)\alpha}-v^{(2)\alpha})]-\mu(t)w=0.$
By mean-value expansions of the nonlinear term and the maximum principle for parabolic equations, one concludes that $w\equiv 0$, yielding uniqueness.
Step 3.6: Conclusion. We have shown that $v_{-}$ and $v_{+}$ are respectively a subsolution and a supersolution of (1.1) with boundary condition (1.2), that they are ordered, and that the initial data lies between them. Invoking the standard subsolution–supersolution existence theorem (see [11], Theorem 1.3 and related results, or [8], Section 3), we conclude that there exists a unique classical solution $v$ of the problem (1.1)–(1.2)–(1.3) satisfying the pointwise estimate (2.16). This completes the proof. ∎
In the following section, we present the derivation of the model from fundamental conservation principles, connecting it to the classical framework of Gurtin and MacCamy.
3. Derivation of the Model via Conservation Laws
The nonlinear diffusion equation (1.1) can be derived from the fundamental conservation law:
$\frac{\partial v}{\partial t}+\nabla\cdot\mathbf{J}=f(v,t),$ (3.1)
where $v(x,t)$ represents the population density, $\mathbf{J}$ is the population flux, and $f(v,t)$ is the net growth rate. We consider a constitutive function
$\varphi:[0,\infty)\rightarrow[0,\infty),$
which is differentiable in $v$ and satisfies the conditions:
$\varphi^{\prime}(0)=0,\quad\varphi^{\prime}(v)>0\quad\text{for }v>0.$
In our model, the population flux $\mathbf{J}$ is governed by this potential $\varphi$, following a nonlinear Fick’s law:
$\mathbf{J}=-\nabla\varphi(v)=-\varphi^{\prime}(v)\nabla v.$ (3.2)
This indicates that the dispersal is driven by the gradient of the potential $\varphi(v)$. Substituting (3.2) into (3.1) yields the diffusion term $\nabla\cdot(\varphi^{\prime}(v)\nabla v)$.
This flux-driven dynamic can be intimately connected to a variational principle. The Euler–Lagrange conditions associated with the functional are inspired by the population flux $\mathbf{J}$, suggesting that the system evolves to minimize an associated energy. We consider:
$E(v)=\int_{\Omega}\left[\frac{1}{2}\Bigl(\varphi^{\prime}(v)\Bigr)^{2}+\frac{\mu\left(t\right)}{2}v^{2}\right]dx.$
Minimization of the energy $E(v)$, by using a flow in the positive direction of the functional derivative, leads to:
$\frac{\partial v}{\partial t}=\frac{\delta E}{\delta v}=\nabla\cdot\Bigl(\varphi^{\prime}(v)\nabla v\Bigr)+\mu\left(t\right)v.$
Here, the variational derivation recovers the exact form of the flux term derived from conservation laws. We observe that the diffusive term $\Delta\varphi(v)$ can be expressed in the form:
$\Delta\varphi(v)=\nabla\cdot\Bigl(\varphi^{\prime}(v)\nabla v\Bigr).$
By choosing the specific function $\varphi(v)=v^{\alpha}$, the equation (1.1) is obtained after rearrangement. The interpretations are as follows: the diffusive term $\Delta v^{\alpha}$ describes the dispersal of individuals in a habitat (representing the way in which the population spreads out in space), while the term $\mu\left(t\right)v$ represents the effect of population growth or mortality, modeling the birth and death processes. The boundary conditions (1.2) and the initial condition (1.3) complement the model by ensuring that the initial population distribution is predefined and that there is no flux of individuals across the domain’s boundary.
4. Conclusions
In this work, we have established a comprehensive theoretical and computational framework for the nonlinear diffusion equation of Gurtin–MacCamy type with time-dependent growth rates. Our main contributions can be summarized as follows:
  1. 1.
    Theorem 1.1: We proved the existence and uniqueness of separable classical solutions of the form
    $v(x,t)=S(t)u^{1/\alpha}(x),$
    where the spatial profile $u(x)$ satisfies a sublinear elliptic equation and the temporal function $S(t)$ is governed by a Bernoulli-type ODE. This separation-of-variables approach provides explicit analytical expressions that are directly implementable.
  2. 2.
    Theorem 1.3: We extended the existence and uniqueness theory to the general non-separable case using the monotone iteration method based on subsolution–supersolution techniques. This result demonstrates that solutions remain bounded within the ordered interval $v_{-},v_{+}]$ for arbitrary initial data satisfying the compatibility conditions.
  3. 3.
    Theorem 1.2: We established a more general comparison principle, showing that if the initial population $v_{0}$ is smaller than $w_{0}$ and the growth rate $\mu_{1}(t)$ is smaller than $\mu_{2}(t)$, then the population level $v(x,t)$ remains bounded by $w(x,t)$ for all subsequent times.
  4. 4.
    Numerical Validation: Python implementations of both theorems verify the consistency between the analytical separable solution and the general iterative approach. The simulations demonstrate that the iterative scheme converges to the exact separable form when applicable.
The parameter sensitivity analysis reveals that: the nonlinear exponent $\alpha$ governs the diffusion rate and solution regularity; the time-dependent growth rate $\mu(t)$ controls the temporal dynamics and long-time behavior; and the initial parameter $\gamma$ determines the amplitude scaling of the solution.
The proposed framework has broad applicability beyond population dynamics, including thermal diffusion in non-homogeneous materials and the modeling of bacterial contamination in healthcare environments. Future research directions include: extension to higher-dimensional domains with complex geometries; incorporation of spatially varying diffusion coefficients; and analysis of equations with nonlinear boundary conditions.
Details of the full Python implementations are provided in Appendices AC, enabling reproducibility and practical application of the theoretical results.
Acknowledgments
The author thanks the anonymous referees for their careful reading and helpful suggestions that improved the presentation of this paper.
Funding
This research received no external funding.
Data Availability
Data sharing is not applicable since the article is purely theoretical in nature.
Declarations
Conflicts of Interests The author declares no conflict of interest.
Appendix A Python Implementation of Theorem 1.1
The Python script implementing the construction of the separable solution
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
as established in Theorem 1.1, is publicly available at: https://github.com/coveidragos/PorousMediumEquation/blob/main/Theorem_First.py.
Appendix B Python Implementation of Theorem 1.3
The Python code available at: https://github.com/coveidragos/PorousMediumEquation/blob/main/Theorem_Second.py implements Theorem 1.3 using the subsolution–supersolution framework. In the special case where the solution admits a separable form
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
as in Theorem 1.1, this implementation confirms that both theorems lead to identical numerical outcomes.
Appendix C Python Implementation of Remark 1.4
The Python script available at: https://github.com/coveidragos/PorousMediumEquation/blob/main/Theorem_Second.py extends the implementation from Section B to illustrate Theorem 1.3 under the general initial condition proposed in Remark 1.4. In particular, the parabolic solver is initialized with a weighted average of the sub- and supersolutions,
$v_{0}(x)=\frac{7}{8}v_{-}(x,0)+\frac{1}{8}v_{+}(x,0),$
thereby demonstrating that the monotone iteration scheme robustly captures the existence and uniqueness of solutions for arbitrary initial data within the ordered interval $v_{-},v_{+}]$.
Appendix D Image Denoising Application and Comparison with Perona–Malik
From an application perspective, the nonlinear diffusion models studied in this work, particularly the Gurtin–MacCamy (GM) type equations, find significant utility in the field of digital image processing. Specifically, the power-law diffusion mechanism governed by (1.1) can be effectively employed for image denoising. While classical linear diffusion (the heat equation) tends to blur important structural features like edges, the nonlinear approach preserves sharp transitions while smoothing homogeneous regions.
In this section, we demonstrate that our framework is successfully applicable to denoising and yields results comparable to the established Perona–Malik (PM) [12, 13] anisotropic diffusion model. The performance of our model, along with the metrics for the noisy image and the Perona–Malik results, is summarized in Table 1 for two different noise levels ($\sigma=0.1$ and $\sigma=0.18$).
Table 1: Comparative performance metrics for image denoising on lenna.png.
Noise Level Metric Noisy Image Gurtin–MacCamy Perona–Malik
$\sigma=0.1$ MSE 0.009564 0.002459 0.002529
PSNR (dB) 20.19 26.09 25.97
SSIM 0.2960 0.6583 0.6355
$\sigma=0.18$ MSE 0.028305 0.004723 0.015066
PSNR (dB) 15.48 23.26 18.22
SSIM 0.1533 0.5249 0.2226
These results indicate that while both models perform similarly at low noise levels, the Gurtin–MacCamy approach preserves image structure significantly better as the noise increases. The following Python code available at: https://github.com/coveidragos/PorousMediumEquation/blob/main/Gurtin_Perona_Image.py performs the comparative analysis, showing that the Gurtin–MacCamy model, with properly tuned parameters and a time-dependent growth rate $\mu(t)$, robustly recovers image quality from Gaussian noise.