### v3

# A Bernoulli-Type Localization Principle for Nonlinear Diffusion Equations: Structural Description of Solutions and First Application to Image Processing

Abstract
This paper introduces a genuinely new approach to the analysis of nonlinear diffusion equations of porous medium type with time-dependent growth rates: the localization of solutions through an associated Bernoulli-type ordinary differential equation. This principle—which we term the Bernoulli localization principle—provides, for the first time, a complete structural description of solutions to the initial-boundary-value problem by explicitly decoupling the temporal dynamics, governed by a Bernoulli-type ODE whose solution is available in closed form, from the spatial profile, determined by a time-independent sublinear elliptic equation of Brezis–Oswald type. The resulting decomposition changes the way this class of partial differential equations has been viewed: rather than analyzing the full parabolic PDE by purely qualitative methods, one solves a scalar ODE and a stationary elliptic problem, and their combination yields a complete, explicit structural description of every solution. We establish three main results: (i) existence and uniqueness of separable classical solutions via the transformation of the parabolic problem into an elliptic framework coupled with the Bernoulli ODE (Theorem 1.1); (ii) a comparison principle showing that the initial ordering of data and growth rates is preserved throughout evolution (Theorem 1.2); and (iii) a general existence and uniqueness result for non-separable solutions via monotone iteration based on the subsolution–supersolution technique, with explicit localization bounds constructed from the Bernoulli ODE (Theorem 1.3). To demonstrate the applied character of this framework, we present what is, to our knowledge, the first application of a porous-medium-type diffusion equation to digital image processing, specifically image denoising, where the model achieves results that match or exceed those of the established Perona–Malik anisotropic diffusion filter. A complete and transparent description of the Python implementation is provided. The framework is also relevant to other applied domains—physics (thermal diffusion in porous media), biology (population dynamics, tumor growth), economics (diffusion of innovations, wealth distribution), and engineering (groundwater flow, materials processing)—via concrete models and verifiable predictions.
Keywords: Nonlinear diffusion; Bernoulli localization principle; Structural description of solutions; Image denoising; Porous medium equation; Gurtin–MacCamy model; Existence and uniqueness; Monotone iteration; Perona–Malik comparison; Cross-disciplinary applications.
MSC 2020: 35K55; 35K57; 35J20; 35B40; 35D30; 92D25; 94A08.
1. Introduction
1.1. Motivation and the classical viewpoint
The theory of nonlinear diffusion equations has been a central pillar of mathematical analysis and applied mathematics for over half a century. Equations of the form $\partial_{t}v=\Delta v^{\alpha}$ with $\alpha>1$—the so-called porous medium equations—arise naturally in contexts ranging from the flow of gases through porous rock formations to the dispersal of biological populations in heterogeneous environments [24, 16, 2]. When augmented by a reaction or growth term, these equations become indispensable in modeling processes where diffusion and proliferation interact: thermal conduction in materials with temperature-dependent conductivity [12], density-dependent dispersal of biological species [16, 18, 19], groundwater flow through soils with nonlinear permeability [3], and even the propagation of innovations through social networks [23].
Despite decades of intensive study, the methods available for understanding the structure of solutions to these equations have remained largely indirect. Comparison principles, energy estimates, regularity theory, and asymptotic analysis provide existence, uniqueness, and qualitative information, but they typically do not furnish an explicit structural description of the solution. The celebrated work of Brezis and Oswald [5] established the existence and uniqueness of solutions to sublinear elliptic problems, while Brezis and Kamin [4] extended these results to unbounded domains; the monograph of Vázquez [24] gave a comprehensive treatment of existence, uniqueness, regularity, and asymptotics for the porous medium equation; and the foundational framework of Gurtin and MacCamy [16] provided the biological underpinning of density-dependent diffusion. In all these approaches, however, the solution is characterized implicitly: it exists, it is unique, it has certain qualitative properties—but its internal structure remains opaque.
1.2. A new idea: the Bernoulli localization principle
In this paper we introduce a genuinely new approach that fundamentally changes this picture. The central observation is disarmingly simple yet far-reaching: the solutions of the nonlinear parabolic problem can be localized—that is, explicitly described, bounded, and structurally characterized—by means of an associated Bernoulli-type ordinary differential equation. This localization principle rests on a decomposition of the solution into a temporal factor $S(t)$, governed by a scalar Bernoulli ODE with a closed-form solution, and a spatial profile $u^{1/\alpha}(x)$, determined by a stationary elliptic equation. The combination of these two factors provides a complete structural description of the solution that is unprecedented for this class of equations.
To be precise, we study the nonlinear diffusion equation
$\frac{\partial v}{\partial t}(x,t)-\Delta v^{\alpha}(x,t)=\mu\left(t\right)v(x,t),\quad\text{in }\Omega\times(0,T),$ (1.1)
where $\alpha>1$, $T>0$ (possibly infinite), $\mu:\left[0,\infty\right)\rightarrow\left(0,\infty\right)$ is a continuous time-dependent growth rate (additionally bounded whenever $T$ is infinite), and $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded smooth domain with hostile boundaries ($\partial\Omega$). Problem (1.1) is supplemented with boundary and initial conditions:
$v(x,t)$ $=0,\qquad\text{on }\partial\Omega\times(0,T),$ (1.2)
$v(x,0)$ $=v_{0}(x),\qquad\text{on }\Omega.$ (1.3)
Here, $v(x,t)$ represents the evolving population density (or, analogously, temperature, concentration, or image intensity), $\Delta v^{\alpha}(x,t)$ embodies a power-law diffusion process, $\alpha$ governs the nonlinearity of diffusion, $\mu(t)$ represents the time-dependent growth or decay rate, $T$ defines the observation time frame, the boundary condition (1.2) ensures the solution vanishes at the boundary of the domain, and the initial condition (1.3) specifies the initial state.
The key methodological insight of our study is that a separation-of-variables ansatz, combined with the identification of the temporal factor as the solution of a Bernoulli-type ODE, yields not merely an existence result but an explicit structural formula for the solution. Specifically, we show that the problem (1.1)–(1.2)–(1.3) admits solutions of the form
$v_{\gamma}\left(x,t\right)=S\left(t\right)u^{\frac{1}{\alpha}}\left(x\right),\quad S\left(t\right)=\left[e^{\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\left(\gamma^{1-\alpha}+(\alpha-1)\int_{0}^{t}e^{-\left(1-\alpha\right)\int_{0}^{\tau}\mu\left(s\right)ds}d\tau\right)\right]^{\frac{1}{1-\alpha}},$ (1.4)
where $u\left(x\right)$ is the unique positive solution of the associated elliptic problem
$\left\{\begin{array}[]{ccc}-\Delta u\left(x\right)=u^{\frac{1}{\alpha}}\left(x\right)&in&\Omega,\\ u\left(x\right)>0&in&\Omega,\\ u\left(x\right)=0&on&\partial\Omega,\end{array}\right$. (1.5)
and the initial condition is satisfied as
$v_{0}(x)=v_{\gamma}\left(x,0\right)=S\left(0\right)u^{\frac{1}{\alpha}}\left(x\right),$ (1.6)
with $S\left(0\right)=\gamma\in\left(0,\infty\right)$ given.
This representation is not merely a calculational convenience; it reveals the structural anatomy of the solution. The temporal dynamics are entirely governed by a scalar Bernoulli-type ODE, while the spatial profile is determined once and for all by a stationary sublinear elliptic problem. This decomposition has several profound consequences: (a) the long-time behavior is immediately readable from the asymptotics of $S(t)$; (b) parameter sensitivity is transparent, since $\alpha$ affects both $S$ and $u$ in explicitly traceable ways; and (c) numerical implementation reduces to solving an elliptic equation (a mature, well-understood task) plus evaluating the closed-form expression for $S(t)$.
1.3. Literature and context
It is well known that problem (1.1), together with the boundary conditions (1.2)–(1.3), is typically analyzed within the theoretical framework introduced by Gurtin and MacCamy [16], which provides a foundational approach to diffusion processes with density-dependent dispersal. Subsequent research has extended this setting in various directions; for instance, when $\mu\left(t\right)$ is a constant and $\Omega=\mathbb{R}^{N}$, the problem has been studied in detail by He [15]. When $\mu\left(t\right)=0$, the equation (1.1) reduces to the classical porous medium equation,
$\frac{\partial v}{\partial t}(x,t)=\Delta v^{\alpha}(x,t),\quad\text{in }\Omega\times(0,T),$
well known for its degenerate parabolic structure, finite-speed propagation, and rich regularity theory; see [17, 24, 25] for detailed treatments.
In biological systems, nonlinear diffusion equations capture the complex dynamics of microbial populations where growth and dispersal interact in nontrivial ways [16, 19, 18]. Similar models describe thermal processes in non-homogeneous materials, phase transitions, and the evolution of microstructural defects [24, 12]. The foundational work of Gurtin and MacCamy [16] introduced the density-dependent diffusion model for biological populations, where the diffusion coefficient depends on the population density itself, leading to equations of the form $v_{t}=\Delta v^{\alpha}$ with $\alpha>1$ (slow diffusion). The monograph by Vázquez [24] provides a comprehensive treatment of the mathematical theory, while the work of Aronson and Peletier [1] established fundamental results on the large-time behavior of solutions in bounded domains.
The subsolution–supersolution method, systematically developed by Pao [20] for parabolic systems, provides a powerful framework for establishing existence and uniqueness in the general non-separable case. Recent progress on population models with nonlinear survival structures has been reported in [7] and [6]. These contributions, while significant, characterize solutions qualitatively. What has been missing—and what this paper provides—is a principle that makes the solution structure explicit.
1.4. Broader applicability
The broader applicability of the Bernoulli localization framework extends well beyond biological dynamics and encompasses several scientific and engineering disciplines.
In physics, equations of the form (1.1) describe thermal diffusion in porous media where the thermal conductivity is a power law of the temperature [2]. The localization principle predicts explicit upper and lower bounds on the temperature field at any time, given the initial temperature profile.
In biology, density-dependent dispersal models such as (1.1) arise naturally in ecology [19, 18, 13], where $v(x,t)$ represents the density of a species and $\mu(t)$ captures time-varying birth-death dynamics. The structural decomposition allows explicit computation of maximum and minimum population densities at any time $t$.
In economics and social sciences, nonlinear diffusion models capture the spread of innovations through a population [23], where $v(x,t)$ represents adoption density and the power-law diffusion reflects peer-influence effects. The localization principle yields verifiable bounds on market penetration as a function of time.
In engineering, the porous medium equation governs groundwater flow [3], oil recovery from reservoirs, and materials processing (e.g., sintering). In each of these contexts, the Bernoulli localization principle provides quantitative predictions that can be tested against field data or laboratory measurements.
We elaborate on these cross-disciplinary connections in Section 5.
1.5. Main contributions and novelty
In summary, this paper establishes three key theoretical results complemented by a first application to image processing:
  • Theorem 1.1 (Bernoulli localization for separable solutions): An existence and uniqueness result for separable solutions, obtained by transforming the parabolic problem into an elliptic equation coupled with a Bernoulli-type ordinary differential equation. The solution is given in closed form by (1.4).
  • Theorem 1.2 (Comparison principle): A comparison result with respect to the initial conditions and growth rates, showing that the population distribution preserves its initial ordering throughout evolution.
  • Theorem 1.3 (Localization bounds for general solutions): A general existence and uniqueness result for non-separable solutions, established via the monotone iteration method based on subsolution–supersolution techniques, where the sub- and supersolutions are explicitly constructed from the Bernoulli ODE, thereby providing concrete localization bounds.
  • Image denoising application: The first use of a porous-medium-type diffusion equation for image denoising, with a transparent line-by-line explanation of the Python implementation, demonstrating performance comparable to or exceeding the classical Perona–Malik filter.
The novelty of this work is threefold. First, the Bernoulli localization principle itself is new: it provides a structural decomposition that has not been available for this class of equations. Second, the explicit localization bounds in Theorem 1.3—constructed from the temporal factor $S(t)$ and the elliptic sub- and supersolutions—go beyond classical qualitative comparison principles by furnishing quantitatively sharp barriers. Third, the application to image processing is new: the Gurtin–MacCamy diffusion model has not previously been used for image denoising, and our results show that it outperforms the classical Perona–Malik filter, especially at higher noise levels.
1.6. Statement of main results
We now present the main results in their abstract formulation.
Theorem 1.1.
Let $\Omega\subset\mathbb{R}^{N}$ be a bounded, arcwise-connected open set with a compact closure and boundary $\partial\Omega$ of class $C^{3}$. Under these conditions, the nonlinear diffusion problem (1.1)–(1.2) admits a unique solution $v_{\gamma}(x,t)$, given by (1.4), where $u\in C^{2}(\Omega)\cap C^{1}(\overline{\Omega})$ exists uniquely, and the initial boundary condition (1.3) is given by (1.6).
Our methodology for proving Theorem 1.1 departs from classical strategies, such as those found in [4, 16, 25], by harnessing a transformation that simplifies the analysis even under challenging regimes. This approach not only strengthens the theoretical framework but also paves the way for robust numerical implementations.
Theorem 1.2.
Let $v_{0},w_{0}\in C(\overline{\Omega})$ be two initial conditions such that $v_{0}(x)\leq w_{0}(x)$ for all $x\in\overline{\Omega}$. Assume that $\mu_{1}\left(t\right)$ and $\mu_{2}\left(t\right)$ are continuous and bounded functions such that $\mu_{1}\left(t\right)\leq\mu_{2}\left(t\right)$. If $v(x,t)$ and $w(x,t)$ are the unique classical solutions to problem (1.1)–(1.3) corresponding to $(v_{0},\mu_{1})$ and $(w_{0},\mu_{2})$, respectively, then
$v(x,t)\leq w(x,t)\quad\text{for all }\left(x,t\right)\in\overline{\Omega}\times[0,T).$ (1.7)
Intuitively, this result shows that if the initial population $v_{0}$ is smaller than $w_{0}$ and the growth rate $\mu_{1}(t)$ is smaller than $\mu_{2}(t)$, then solution $v$ stays below $w$ during the entire observation period.
Theorem 1.1 and Theorem 1.2 lead to a new result that is presented here for the first time.
Theorem 1.3.
Let $\Omega\subset\mathbb{R}^{N}$ be a bounded, arcwise-connected open set with a compact closure and boundary $\partial\Omega$ of class $C^{3}$. Assume that $u_{-},u_{+}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ satisfy
$-\Delta u_{-}$ $\leq u_{-}^{1/\alpha}\qquad\text{in }\Omega,$ $u_{-}>0\text{ in }\Omega,$ $u_{-}=0\text{ on }\partial\Omega,$ (1.8)
$-\Delta u_{+}$ $\geq u_{+}^{1/\alpha}\qquad\text{in }\Omega,$ $u_{+}>0\text{ in }\Omega,$ $u_{+}=0\text{ on }\partial\Omega,$ (1.9)
and $u_{-}\leq u_{+}$ in $\overline{\Omega}$. Fix $\gamma>0$ and let $S$ be the unique solution of
$S^{\prime}(t)+S^{\alpha}(t)\,=\mu\left(t\right)S\left(t\right),\qquad S(0)=\gamma,$ (1.10)
as given by Equation (1.4). We define $v_{-}(x,t)=S(t)u_{-}^{1/\alpha}(x)$ and $v_{+}(x,t)=S(t)u_{+}^{1/\alpha}(x)$. If $v_{0}\in C(\overline{\Omega})$ satisfies $v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)$ in $\overline{\Omega}$, then there exists a unique classical solution $v(x,t)$ to (1.1)-(1.3), such that $v_{-}(x,t)\leq v(x,t)\leq v_{+}(x,t)$ for all $\left(x,t\right)\in\overline{\Omega}\times\left[0,T\right)$.
A careful examination of Theorems 1.1 and 1.3 and their proofs shows that both results play an essential role, even though Theorem 1.1 is a particular case of Theorem 1.3. Theorem 1.1 remains significant because its proof lends itself to straightforward implementation in a programming language through simplified numerical procedures, whereas Theorem 1.3 applies to a more general situation in which the solution is not necessarily separable. As an illustration, we present a well-known result that highlights the distinct roles played by Theorems 1.1 and 1.3.
Remark 1.4.
There always exists a continuous function
$u_{-}^{+}:\overline{\Omega}\rightarrow[0,\infty)$
such that
$u_{-}(x)\leq u_{-}^{+}(x)\leq u_{+}(x)\quad\text{for all }x\in\overline{\Omega}.$ (1.11)
For instance, for any continuous function $\theta:\overline{\Omega}\rightarrow[0,1]$, the function
$u_{-}^{+}(x):=(1-\theta(x))\,u_{-}(x)+\theta(x)\,u_{+}(x)$
satisfies Eq. (1.11).
Therefore, Theorem 1.2 strengthens both the first (Theorem 1.1) and third result (Theorem 1.3). First, it plays a key role in the proof of Theorem 1.3. Second, when combined with Theorem 1.1, it enables the construction of a subsolution and supersolution pair, which in turn makes the implementation of the results in a computer language feasible.
1.7. Organization of the paper
The remainder of this paper is organized as follows. Section 2 presents the Bernoulli localization principle in detail, explaining why it constitutes a new structural framework for this class of equations. Section 3 provides complete proofs of Theorems 1.11.3. In Section 4, we derive the model from conservation laws and energy considerations, following the framework of Gurtin and MacCamy. Section 5 discusses cross-disciplinary relevance, presenting concrete models and verifiable predictions for physics, biology, economics, and engineering. Section 6 provides the complete Python implementation for image denoising and presents a detailed comparison with the Perona–Malik filter, including a transparent explanation of every component of the code. Finally, Section 7 presents conclusions and future research directions.
2. The Bernoulli Localization Principle: A New Structural Framework
In this section we describe in detail the central conceptual contribution of this paper: the Bernoulli localization principle. We explain what it asserts, why it is new, and how it changes the way one can think about the solutions of porous-medium-type diffusion equations with time-dependent growth.
2.1. The principle
Definition 2.1 (Bernoulli localization).
We say that the solutions of the parabolic problem (1.1)–(1.2)–(1.3) are localized by the Bernoulli ODE if there exist functions $v_{-}(x,t)$ and $v_{+}(x,t)$ of the product form
$v_{\pm}(x,t)=S(t)\,u_{\pm}^{1/\alpha}(x),$ (2.1)
where $S(t)$ is the unique positive solution of the Bernoulli-type ODE
$S^{\prime}(t)+S^{\alpha}(t)=\mu(t)\,S(t),\qquad S(0)=\gamma>0,$ (2.2)
and $_{-}, u+u_{+}$ are, respectively, a subsolution and a supersolution of the stationary elliptic problem (1.5) with $u_{-}\leq u_{+}$ in $\overline{\Omega}$, such that every classical solution $v$ of (1.1)–(1.2)–(1.3) with initial data satisfying $v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)$ is trapped between these barriers:
$v_{-}(x,t)\leq v(x,t)\leq v_{+}(x,t)\quad\text{for all }(x,t)\in\overline{\Omega}\times[0,T).$ (2.3)
The power of this definition lies in the explicit, factored form of the barriers. The temporal factor $S(t)$ is universal: it depends only on $\alpha$, $\mu(\cdot)$, and the initial parameter $\gamma$, not on the spatial domain $\Omega$ or the spatial profile $u$. It is given in closed form by Equation (1.4). The spatial factors $u_{\pm}^{1/\alpha}$ are determined by the stationary elliptic problem on $\Omega$—a problem for which existence, uniqueness, and regularity are classical [5, 14].
2.2. Why the principle is new
The novelty of the Bernoulli localization principle can be appreciated by contrasting it with the approaches that have dominated the literature.
  1. 1.
    Classical comparison principles [20, 24] establish that if a subsolution lies below a supersolution at $t=0$ and on the parabolic boundary, then the same ordering persists for all time. These principles are extremely powerful but remain qualitative: they do not specify the functional form of the sub- and supersolutions, nor do they relate the temporal and spatial dependences explicitly. The Bernoulli localization principle, by contrast, produces barriers of a specific, analytically tractable product form, thereby converting a qualitative ordering into a quantitative structural description.
  2. 2.
    Energy and variational methods [24, 12] establish existence via minimization of suitable energy functionals. These methods yield weak solutions and, through regularity bootstrapping, classical solutions. However, the solutions so obtained remain implicit; the energy functional tells us that a minimizer exists, but not what it looks like. The Bernoulli localization principle, in contrast, tells us precisely what the solution looks like: it is the product of a temporal Bernoulli factor and a spatial elliptic profile.
  3. 3.
    Self-similar and asymptotic analyses [1, 25] provide information about the long-time behavior of solutions, typically showing convergence to a self-similar profile or a steady state. The Bernoulli localization principle subsumes these results: the long-time behavior of $S(t)$ is immediately readable from the explicit formula (1.4), yielding the asymptotic profile without further analysis.
2.3. Structural consequences
The Bernoulli localization principle has several immediate structural consequences.
Complete structural description. Every solution of (1.1)–(1.2)–(1.3) lying in the ordered interval $v_{-},v_{+}]$ is sandwiched between two explicitly known functions. In the separable case, the solution is such a product, and the description is exact rather than merely bounding.
Temporal–spatial decoupling. The full parabolic PDE is replaced by two independent, simpler problems: a scalar Bernoulli ODE for the temporal factor and a stationary elliptic PDE for the spatial profile. Each can be analyzed and solved independently, and their combination yields the solution of the original problem.
Explicit parameter dependence. The dependence of the solution on the parameters $\alpha$, $\gamma$, $\mu(\cdot)$, and $\Omega$ is transparently readable from the formula (1.4) and the elliptic problem (1.5). This facilitates sensitivity analysis, parameter fitting in applications, and the design of numerical experiments.
Algorithmic accessibility. The localization principle translates directly into an algorithm: (i) solve the elliptic problem (1.5) by a standard numerical method (finite elements, finite differences, or the iterative scheme in Theorem 1.1); (ii) evaluate $S(t)$ from the closed-form expression (1.4); (iii) form the product $v(x,t)=S(t)\,u^{1/\alpha}(x)$. No parabolic solver is needed for the separable case.
3. Proof of the Main Results
In the following proof of Theorem 1.1, we apply the successive approximation method, because our goal is to implement the result using a computer language.
Proof of Theorem 1.1.
We first prove the existence of solution to the problem (1.5). Let $\phi_{1}\in C^{2}(\overline{\Omega})$ be the first positive eigenfunction corresponding to the first eigenvalue $\lambda_{1}$ of the problem
$\left\{\begin{array}[]{lll}-\Delta\phi(x)=\lambda\phi(x),&in&\Omega,\\ \phi(x)=0,&on&\partial\Omega,\end{array}\right$.
and $w\in C^{2}(\overline{\Omega})$ be the unique solution of the problem
$\left\{\begin{array}[]{ccc}-\Delta w(x)=1,&in&\Omega,\\ w(x)=0,&on&\partial\Omega.\end{array}\right$.
We choose
$\sigma=\frac{1}{\lambda_{1}^{\frac{\alpha}{\alpha-1}}\max_{x\in\overline{\Omega}}\phi_{1}\left(x\right)}\text{ and }c=\max\left\{\sigma\lambda_{1}\max_{x\in\overline{\Omega}}\phi_{1},\left(\max_{x\in\overline{\Omega}}w\left(x\right)\right)^{\frac{1}{\alpha-1}}\right\}.$
Then, the function $\underline{u}(x)=\sigma\phi_{1}$ is a subsolution of (1.5) whereas $\overline{u}\left(x\right)=cw(x)$ is a supersolution. Indeed,
$-\Delta\underline{u}(x)-\left(\underline{u}(x)\right)^{\frac{1}{\alpha}}$ $=\sigma\lambda_{1}\phi_{1}\left(x\right)-\left(\sigma\phi_{1}\right)^{\frac{1}{\alpha}}$
$=\sigma\lambda_{1}\phi_{1}\left(x\right)-\sigma^{\frac{1}{\alpha}}\left(\phi_{1}\left(x\right)\right)^{\frac{1}{\alpha}}$
$=\sigma\left(\lambda_{1}\phi_{1}\left(x\right)\right)^{\frac{1}{\alpha}}\left[\left(\lambda_{1}\phi_{1}\left(x\right)\right)^{\frac{\alpha-1}{\alpha}}-\frac{1}{\sigma}\left(\frac{\sigma}{\lambda_{1}}\right)^{\frac{1}{\alpha}}\right]$
$\leq\sigma\left(\lambda_{1}\phi_{1}\left(x\right)\right)^{\frac{1}{\alpha}}\left[\left(\lambda_{1}\max_{x\in\overline{\Omega}}\phi_{1}\left(x\right)\right)^{\frac{\alpha-1}{\alpha}}-\frac{1}{\sigma}\left(\frac{\sigma}{\lambda_{1}}\right)^{\frac{1}{\alpha}}\right]=0$
and
$-\Delta\overline{u}(x)-\left(\overline{u}\left(x\right)\right)^{\frac{1}{\alpha}}$ $=c-\left(cw\left(x\right)\right)^{\frac{1}{\alpha}}\geq c-\left(c\max_{x\in\overline{\Omega}}w\left(x\right)\right)^{\frac{1}{\alpha}}$
$=c^{\frac{1}{\alpha}}\left[c^{\frac{\alpha-1}{\alpha}}-\left(\max_{x\in\overline{\Omega}}w\left(x\right)\right)^{\frac{1}{\alpha}}\right]\geq 0.$
Now, since
$\left\{\begin{array}[]{lll}-\Delta[\overline{u}(x)-\underline{u}(x)]=c-\sigma\lambda_{1}\phi_{1}\geq c-\sigma\lambda_{1}\max_{x\in\overline{\Omega}}\phi_{1}\geq 0&in&\Omega\\ \overline{u}(x)-\underline{u}(x)=0&on&\partial\Omega\end{array}\right$.
it follows from the maximum principle that $\underline{u}(x)\leq\overline{u}(x)$, for all $x\in\overline{\Omega}$.
Next, we define the sequence $\{u^{k}\}_{k\geq 0}$ as
$u^{0}(x)=u_{0}(x),\text{ }-\Delta u^{k}\left(x\right)=\left(u^{k-1}\left(x\right)\right)^{\frac{1}{\alpha}}\text{ for }k=1,2,...\text{ and }x\in\Omega$ (3.1)
where
$u_{0}(x)=\underline{u}(x)\quad\forall\,x\in\overline{\Omega},$
with $\underline{u}(x)$ being a subsolution of the problem
$\Delta u(x)=u(x)^{\frac{1}{\alpha}},\quad u(x)>0\text{ in }\Omega,\qquad u(x)=0\text{ on }\partial\Omega.$ (3.2)
Using the Brezis-Oswald inequality [5, Theorems 1 and 2], we derive
$\underline{u}(x)\leq u^{0}(x)\leq u^{1}(x)\leq u^{2}(x)\leq\cdots\leq\overline{u}(x),\quad\forall\,x\in\overline{\Omega}.$ (3.3)
Because the sequence $\{u^{k}\}_{k\geq 0}$ is monotonically increasing and bounded above (by $\overline{u}(x)$), it follows from the monotone convergence theorem that the pointwise limit
$u(x)=\lim_{k\rightarrow\infty}u^{k}(x),$
exists for all $x\in\overline{\Omega}$. Standard regularity theory, which can be found in [14], ensures that $u(x)\in C^{2}(\Omega)\cap C^{1}(\overline{\Omega})$ and, by construction, satisfies the required conditions for problem (1.5).
$\underline{u}(x)\leq u(x)\leq\overline{u}(x)\quad\forall\,x\in\overline{\Omega}.$
The uniqueness of the solution to problem (1.5) follows again from the results established by Brezis and Oswald [5, Theorems 1 and 2] (see also [1] or [4]).
Now, let us prove the validity of (1.4). Assume that a solution of the form
$v(x,t)=u^{\frac{1}{\alpha}}(x)\,S(t).$ (3.4)
Then we have
$v^{\alpha}(x,t)=u(x)\,S^{\alpha}(t).$
Substitute (3.4) into (1.1) to obtain:
$u^{\frac{1}{\alpha}}(x)\,S^{\prime}(t)-S^{\alpha}(t)\,\Delta u\left(x\right)=\mu\left(t\right)u^{\frac{1}{\alpha}}(x)\,S(t).$ (3.5)
Dividing (3.5) by $u^{\frac{1}{\alpha}}(x)\,$, we have
$S^{\prime}(t)-S^{\alpha}(t)\,\frac{\Delta u\left(x\right)}{u^{\frac{1}{\alpha}}(x)}=\mu\left(t\right)S\left(t\right),$ (3.6)
or, equivalently
$S^{\prime}(t)+S^{\alpha}(t)\,=\mu\left(t\right)S\left(t\right),$ (3.7)
a Bernoulli-type equation. Multiplying both sides of (3.7) by $S^{-\alpha}(t)$ yields
$S^{\prime}(t)S^{-\alpha}(t)\,=\mu\left(t\right)S^{1-\alpha}\left(t\right)-1.$
We now solve this equation by the standard substitution. Define
$z(t)=S(t)^{\,1-\alpha}.$ (3.8)
Then, by differentiating in (3.8), we have
$z^{\prime}(t)=\left(1-\alpha\right)S^{-\alpha}(t)S^{\prime}\left(t\right)\Longrightarrow S^{-\alpha}(t)S^{\prime}\left(t\right)=\frac{z^{\prime}(t)}{1-\alpha}.$
Substitute the expression for $S^{-\alpha}(t)S^{\prime}\left(t\right)$ into the above:
$\frac{z^{\prime}(t)}{1-\alpha}=\mu\left(t\right)z(t)-1.$
This simplifies to
$z^{\prime}(t)=\left(1-\alpha\right)\mu\left(t\right)z(t)+\alpha-1.$ (3.9)
Multiplying both sides of (3.9) by $e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}$ yields
$\left(z\left(t\right)e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\right)^{\prime}=\left(\alpha-1\right)e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds},$
from where
$z\left(t\right)=e^{\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\int\left(\alpha-1\right)e^{-\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}dt.$
But recall that $S(t)^{\,1-\alpha}=z(t)$; hence,
$S\left(t\right)=\left[e^{\left(1-\alpha\right)\int_{0}^{t}\mu\left(s\right)ds}\left(\gamma^{1-\alpha}+(\alpha-1)\int_{0}^{t}e^{-\left(1-\alpha\right)\int_{0}^{\tau}\mu\left(s\right)ds}d\tau\right)\right]^{\frac{1}{1-\alpha}}.$ (3.10)
Replacing (3.10) in (3.4) we obtain the form (1.4). This concludes the proof of Theorem 1.1. ∎
Remark 3.1.
The Python script implementing the construction of the separable solution
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
as established in Theorem 1.1, is publicly available at [8].
Remark 3.2.
The proof of Theorem 1.1 reveals the essential mechanism behind the Bernoulli localization principle: the substitution of the product ansatz (3.4) into the parabolic PDE (1.1) automatically produces the Bernoulli ODE (3.7), because the spatial dependence cancels exactly when $u$ solves the elliptic problem (1.5). This cancellation is not an accident; it is a structural feature of the power-law nonlinearity $v^{\alpha}$, which ensures that $\Delta v^{\alpha}=S^{\alpha}\Delta u$ is proportional to $v$ through the elliptic equation. This structural observation is the conceptual heart of the localization principle.
We then establish a proof of the comparison principle.
Proof of Theorem 1.2.
Let
$\mathcal{L}_{\mu}[u]:=\frac{\partial u}{\partial t}-\Delta u^{\alpha}-\mu(t)u$
be the parabolic operator associated with (1.1). By hypothesis, $v$ and $w$ are unique classical solutions satisfying
$\mathcal{L}_{\mu_{1}}[v]=0\text{ and }\mathcal{L}_{\mu_{2}}[w]=0\text{ in }\Omega\times(0,T).$
We first establish that $v$ serves as a subsolution for the operator $\mathcal{L}_{\mu_{2}}$. Indeed, substituting $v$ into $\mathcal{L}_{\mu_{2}}$ yields:
$\mathcal{L}_{\mu_{2}}[v]=\frac{\partial v}{\partial t}-\Delta v^{\alpha}-\mu_{2}(t)v=\mu_{1}(t)v-\mu_{2}(t)v=(\mu_{1}(t)-\mu_{2}(t))v.$ (3.11)
Because $\mu_{1}(t)\leq\mu_{2}(t)$ for all $t$ and $v(x,t)\geq 0$ in $\Omega\times(0,T)$, it follows that $(\mu_{1}(t)-\mu_{2}(t))v\leq 0$, which implies:
$\frac{\partial v}{\partial t}-\Delta v^{\alpha}\leq\mu_{2}(t)v\quad\text{in }\Omega\times(0,T).$ (3.12)
On the other hand, the solution $w$ satisfies:
$\frac{\partial w}{\partial t}-\Delta w^{\alpha}=\mu_{2}(t)w\quad\text{in }\Omega\times(0,T).$ (3.13)
Furthermore, the initial and boundary conditions satisfy
$v(x,0)=v_{0}(x)\leq w_{0}(x)=w(x,0)\text{ in }\Omega\text{,}$
and
$v(x,t)=0=w(x,t)\text{ on }\partial\Omega\times(0,T).$
To establish the ordering $v\leq w$, we define
z(x,t):=v(x,t)w(x,t)z(x,t):=v(x,t)-w(x,t)
and note that $z$ satisfies:
$\frac{\partial z}{\partial t}-\Delta(v^{\alpha}-w^{\alpha})=\mu_{1}(t)v-\mu_{2}(t)w\leq\mu_{2}(t)(v-w)=\mu_{2}(t)z.$ (3.14)
Using the monotonicity of the mapping $s\mapsto s^{\alpha}$ for $\alpha>1$ and $s\geq 0$, we can express the difference $v^{\alpha}-w^{\alpha}$ as c(x,t)(vw)c(x,t)(v-w) where
$c(x,t)=\int_{0}^{1}\alpha[w+\theta(v-w)]^{\alpha-1}d\theta\geq 0.$
Substituting this into the inequality for $z$, we obtain:
$\frac{\partial z}{\partial t}-\Delta(c(x,t)z)-\mu_{2}(t)z\leq 0.$ (3.15)
Given that $z(x,0)\leq 0$ and $z$ vanishes at the boundary, the application of the maximum principle for porous-medium-type equations (see [24, Chapter 3] or [20, Theorem 2.1]) ensures that $z(x,t)\leq 0$ on $\overline{\Omega}\times[0,T)$. Consequently, $v(x,t)\leq w(x,t)$ for all $(x,t)\in\overline{\Omega}\times[0,T)$, completing the proof. ∎
We conclude this section by establishing the proof of the third main result.
Proof of Theorem 1.3.
First we compute the quantities appearing in the parabolic operator for the functions $v_{\pm}$.
Step 1: Structure of $v_{\gamma}$. Let $u$ be any of $u_{-}$ or $u_{+}$, and define
$v_{\gamma}(x,t):=S(t)\,u^{1/\alpha}(x).$
Then
$v_{\gamma}^{\alpha}(x,t)=\bigl(S(t)\,u^{1/\alpha}(x)\bigr)^{\alpha}=S(t)^{\alpha}\,u(x).$
Hence, using that $S$ depends only on $t$,
$\Delta\bigl(v_{\gamma}^{\alpha}(x,t)\bigr)=\Delta\bigl(S(t)^{\alpha}u(x)\bigr)=S(t)^{\alpha}\,\Delta u(x).$
Moreover,
$\frac{\partial v_{\gamma}}{\partial t}(x,t)=S^{\prime}\left(t\right)u^{1/\alpha}(x).$
Therefore, the left-hand side of the parabolic operator applied to $v_{\gamma}$ is
$\frac{\partial v_{\gamma}}{\partial t}(x,t)-\Delta\bigl(v_{\gamma}^{\alpha}(x,t)\bigr)-\mu(t)\,v_{\gamma}(x,t)$ $=S^{\prime}\left(t\right)u^{1/\alpha}(x)-S(t)^{\alpha}\,\Delta u(x)-\mu(t)\,S(t)\,u^{1/\alpha}(x)$
$=\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u^{1/\alpha}(x)-S(t)^{\alpha}\,\Delta u(x).$
Step 2: Use of the elliptic inequalities.
(a) Subsolution case. Assume $u=u_{-}$ satisfies (1.8), i.e.
$\Delta u_{-}\leq u_{-}^{1/\alpha}\quad\text{in }\Omega.$
This is equivalent to
$\Delta u_{-}\geq-u_{-}^{1/\alpha}\quad\text{in }\Omega.$
Then, for $v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x)$, we have
$\frac{\partial v_{-}}{\partial t}-\Delta\bigl(v_{-}^{\alpha}\bigr)-\mu(t)\,v_{-}$ $=\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{-}^{1/\alpha}-S(t)^{\alpha}\,\Delta u_{-}$
$\leq\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{-}^{1/\alpha}+S(t)^{\alpha}\,u_{-}^{1/\alpha}$
$=\bigl(-S^{\alpha}(t)\bigr)\,u_{-}^{1/\alpha}+S(t)^{\alpha}\,u_{-}^{1/\alpha}=0,$
where we have used (1.10). Therefore,
$\frac{\partial v_{-}}{\partial t}-\Delta\bigl(v_{-}^{\alpha}\bigr)-\mu(t)\,v_{-}\leq 0\quad\text{in }\Omega\times(0,T),$
which shows that $v_{-}$ is a subsolution of (1.1). On the boundary, because $u_{-}=0$ on $\partial\Omega$, we have
$v_{-}(x,t)=S(t)\,0^{1/\alpha}=0$
on $\partial\Omega\times(0,T)$.
(b) Supersolution case. Assume $u=u_{+}$ satisfies (1.9), that is
$\Delta u_{+}\geq u_{+}^{1/\alpha}\quad\text{in }\Omega,$
equivalently
$\Delta u_{+}\leq-u_{+}^{1/\alpha}\quad\text{in }\Omega.$
Then, for $v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x)$, we obtain
$\frac{\partial v_{+}}{\partial t}-\Delta\bigl(v_{+}^{\alpha}\bigr)-\mu(t)\,v_{+}$ $=\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{+}^{1/\alpha}-S(t)^{\alpha}\,\Delta u_{+}$
$\geq\bigl(S^{\prime}(t)-\mu(t)S(t)\bigr)\,u_{+}^{1/\alpha}+S(t)^{\alpha}\,u_{+}^{1/\alpha}$
$=\bigl(-S^{\alpha}(t)\bigr)\,u_{+}^{1/\alpha}+S(t)^{\alpha}\,u_{+}^{1/\alpha}.$
Again, by (1.10), the factor in parentheses vanishes, so
$\frac{\partial v_{+}}{\partial t}-\Delta\bigl(v_{+}^{\alpha}\bigr)-\mu(t)\,v_{+}\geq 0\quad\text{in }\Omega\times(0,T),$
where $v_{+}$ is the supersolution to (1.1). The boundary condition $v_{+}=0$ on $\partial\Omega\times(0,T)$ follows from $u_{+}=0$ on $\partial\Omega$.
Step 3: Ordering of the solution via comparison.
Assume now that $v$ is a classical solution of (1.1) with
$v(x,0)=v_{0}(x)\quad\text{in }\Omega,\qquad v(x,t)=0\quad\text{on }\partial\Omega\times(0,T),$
and that
$v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)\quad\text{for all }x\in\Omega.$
By construction,
$v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x),\qquad v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x),\quad(x,t)\in\overline{\Omega}\times[0,T).$
Therefore the initial data are sandwiched between the sub- and supersolutions at time $t=0$. Moreover, for all $x\in\overline{\Omega}$ yields and $t\in[0,T)$ we have
$v_{-}(x,t)\leq v_{+}(x,t).$
To complete this argument, we invoke classical subsolution–supersolution theory for nonlinear parabolic equations. We proceeded in several steps.
Step 3.1: Initial ordering. By hypothesis, the initial data $v_{0}$ satisfies
$v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)\quad\text{for all }x\in\overline{\Omega}.$
Since
[…]
and because $u_{-}\leq u_{+}$ in $\overline{\Omega}$, this ordering is consistent with the assumptions.
Step 3.2: Boundary compatibility. On the boundary $\partial\Omega\times(0,T)$, both $v_{-}$ and $v_{+}$ vanish:
$v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x)=0,\qquad v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x)=0,$
since $u_{-}=u_{+}=0$ on $\partial\Omega$. This ensures that both the subsolution and supersolution are compatible with the homogeneous Dirichlet boundary condition (1.2).
Step 3.3: Application of comparison principles. The theory developed by Pao [20] (see also [16] for related results) establishes that, under suitable regularity assumptions, if $v_{-}$ and $v_{+}$ are respectively an ordered subsolution and supersolution of the parabolic problem (1.1) with boundary conditions (1.2), and if they satisfy the initial ordering at $t=0$, then there exists a unique classical solution
$v\in C^{2,1}(\Omega\times(0,T))\cap C(\overline{\Omega}\times[0,T)),$
of the problem (1.1)–(1.2)–(1.3) such that
$v_{-}(x,t)\leq v(x,t)\leq v_{+}(x,t)\quad\text{for all }(x,t)\in\overline{\Omega}\times[0,T).$ (3.16)
Step 3.4: Verification of the regularity hypotheses. The regularity assumptions required by the theory are satisfied:
  • (i)
    The domain $\Omega$ is of class $C^{3}$ with compact closure;
  • (ii)
    The functions $u_{-},u_{+}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ are strictly positive in $\Omega$ and vanish on $\partial\Omega$;
  • (iii)
    The function $S\in C^{1}([0,T])$ is strictly positive (since $\gamma>0$ and the ODE (1.10) preserves positivity);
  • (iv)
    The initial data $v_{0}\in C(\overline{\Omega})$ is sandwiched between $v_{-}(\cdot,0)$ and $v_{+}(\cdot,0)$.
These conditions ensure that the nonlinear parabolic operator
$\mathcal{L}[v]:=\frac{\partial v}{\partial t}-\Delta(v^{\alpha})-\mu(t)v$
is well defined and the subsolution–supersolution method applies.
Step 3.4.1: Monotone iterative construction. To make the existence proof constructive and suitable for numerical implementation, we describe the monotone iteration scheme following Pao [20]. This provides an algorithm for approximating the solution $v$ via sequences that converge monotonically from below and above.
(a) Choice of monotonicity constant. We first choose a function $c(x,t)\geq 0$ that is sufficiently large to ensure monotonicity. Specifically, we consider the nonlinearity in (1.1), as follows:
$f(v,t):=\mu(t)v.$
Since $\frac{\partial f}{\partial v}=\mu(t)$, we can take
$c(t):=\max\{0,\,\sup_{s\in[0,t]}|\mu(s)|\},$ (3.17)
where clearly in the case $\mu(s)>0$ we obtain $c(t)=\sup_{s\in[0,t]}\mu(s)$. This ensures that the auxiliary function
$F(v,t):=c(t)\,v+f(v,t)=\bigl(c(t)+\mu(t)\bigr)v$
is nondecreasing in $v$ for $v\in[v_{-},v_{+}]$.
(b) Iterative scheme. We define the two sequences $\{v_{-}^{(m)}\}_{m=0}^{\infty}$ and $\{v_{+}^{(m)}\}_{m=0}^{\infty}$ as follows. Initialize with
$v_{-}^{(0)}(x,t):=v_{-}(x,t),\qquad v_{+}^{(0)}(x,t):=v_{+}(x,t).$ (3.18)
For $m\geq 1$, construct $v_{-}^{(m)}$ and $v_{+}^{(m)}$ by solving the linearized parabolic problem
$\frac{\partial v^{(m)}}{\partial t}-\Delta\bigl((v^{(m)})^{\alpha}\bigr)+c(t)\,v^{(m)}$ $=c(t)\,v^{(m-1)}+\mu(t)\,v^{(m-1)}\qquad\text{in }\Omega\times(0,T),$ (3.19)
$v^{(m)}(x,t)$ $=0\qquad\text{on }\partial\Omega\times(0,T),$
$v^{(m)}(x,0)$ $=v_{0}(x)\qquad\text{in }\Omega.$
Here, the superscript $(m)$ denotes the iteration index, and $v^{(m-1)}$ is the previous iterate.
(c) Monotonicity and convergence. According to the comparison principle for parabolic equations and the choice of $c(t)$, the sequences satisfy
$v_{-}=v_{-}^{(0)}\leq v_{-}^{(1)}\leq v_{-}^{(2)}\leq\cdots\leq v_{-}^{(m)}\leq\cdots\leq v_{+}^{(m)}\leq\cdots\leq v_{+}^{(2)}\leq v_{+}^{(1)}\leq v_{+}^{(0)}=v_{+}.$ (3.20)
That is, the lower sequence $\{v_{-}^{(m)}\}$ monotonically increases and is bounded above by $v_{+}$, whereas the upper sequence $\{v_{+}^{(m)}\}$ monotonically decreases and is bounded below by $v_{-}$. Standard parabolic regularity estimates (see [20], Theorem 2.1) ensure that both sequences converge uniformly on compact subsets of $\overline{\Omega}\times[0,T)$ to limits
$\underline{v}(x,t):=\lim_{m\rightarrow\infty}v_{-}^{(m)}(x,t),\qquad\overline{v}(x,t):=\lim_{m\rightarrow\infty}v_{+}^{(m)}(x,t).$
Moreover, both $\underline{v}$ and $\overline{v}$ are classical solutions of (1.1)–(1.2)–(1.3), and they are the minimal and maximal solutions, in the ordered interval $v_{-},v_{+}]$.
(d) Uniqueness implies convergence to the same limit. If the solution is unique (as established in Step 3.5), then $\underline{v}\equiv\overline{v}\equiv v$, and both sequences converge to the same solution:
$\lim_{m\to\infty}v^{(m)}_{\pm}(x,t)=v(x,t)\quad\text{uniformly on }\overline{\Omega}\times[0,T).$ (3.21)
This constructive approach is particularly well-suited for numerical implementation. At each iteration $m$, one solves a linear parabolic problem (because the right-hand side depends only on the previous iterate $v^{(m-1)}$), which can be discretized using standard finite difference or finite element methods. Monotone convergence (3.20) provides a natural stopping criterion based on the $L^{\infty}$ or $L^{2}$ norm of the difference $\|v^{(m)}-v^{(m-1)}\|$.
Step 3.5: Uniqueness. The uniqueness of the classical solution is derived from the monotonicity of the operator. Specifically, if $v^{(1)}$ and $v^{(2)}$ are two solutions that lie in the interval $v_{-},v_{+}]$, then the difference $w:=v^{(1)}-v^{(2)}$ satisfies
$\frac{\partial w}{\partial t}-\text{div}[\nabla(v^{(1)\alpha}-v^{(2)\alpha})]-\mu(t)w=0.$
By mean-value expansions of the nonlinear term and the maximum principle for parabolic equations, we conclude that $w\equiv 0$, yielding uniqueness.
Step 3.6: Conclusion. We have shown that $v_{-}$ and $v_{+}$ are respectively a subsolution and a supersolution of (1.1) with boundary condition (1.2), that they are ordered, and that the initial data lie between them. Invoking the standard subsolution–supersolution existence theorem (see [20], Theorem 1.3 and related results, or [16], Section 3), we conclude that there exists a unique classical solution $v$ of problem (1.1)–(1.2)–(1.3) satisfying the pointwise estimate (3.16). This completes this proof. ∎
Remark 3.3.
The Python code available at: [9] implements Theorem 1.3 using the subsolution–supersolution framework. In the special case where the solution admits a separable form
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
as in Theorem 1.1, this implementation confirms that both theorems lead to identical numerical outcomes.
Remark 3.4.
The Python script available at: [10] extends the implementation to illustrate Theorem 1.3 under the general initial condition proposed in Remark 1.4. In particular, the parabolic solver is initialized with a weighted average of the sub- and supersolutions,
$v_{0}(x)=\frac{7}{8}v_{-}(x,0)+\frac{1}{8}v_{+}(x,0),$
thereby demonstrating that the monotone iteration scheme robustly captures the existence and uniqueness of the solutions for arbitrary initial data within the ordered interval $v_{-},v_{+}]$.
In the following section, we present the derivation of the model from fundamental conservation principles, connecting it to the classical framework of Gurtin and MacCamy.
4. Derivation of the Model via Conservation Laws
The nonlinear diffusion equation (1.1) can be derived from the fundamental conservation law:
$\frac{\partial v}{\partial t}+\nabla\cdot\mathbf{J}=f(v,t),$ (4.1)
where $v(x,t)$ is the population density, $\mathbf{J}$ is the population flux, and $f(v,t)$ is the net growth rate. The following constitutive function is considered:
$\varphi:[0,\infty)\rightarrow[0,\infty),$
which is differentiable in $v$ and satisfies the conditions:
$\varphi^{\prime}(0)=0,\quad\varphi^{\prime}(v)>0\quad\text{for }v>0.$
In our model, the population flux $\mathbf{J}$ is governed by the potential $\varphi$, following a nonlinear Fick’s law:
$\mathbf{J}=-\nabla\varphi(v)=-\varphi^{\prime}(v)\nabla v.$ (4.2)
This indicates that dispersal is driven by the gradient of the potential $\varphi(v)$. Substituting (4.2) into (4.1) yields the diffusion term $\nabla\cdot(\varphi^{\prime}(v)\nabla v)$.
This flux-driven dynamic can be connected to a variational principle. The Euler–Lagrange conditions associated with the functional are inspired by the population flux $\mathbf{J}$, which suggest that the system evolves to minimize the associated energy. We consider:
$E(v)=\int_{\Omega}\left[\frac{1}{2}\Bigl(\varphi^{\prime}(v)\Bigr)^{2}+\frac{\mu\left(t\right)}{2}v^{2}\right]dx.$
Minimization of the energy $E(v)$, by using a flow in the positive direction of the functional derivative, leads to:
$\frac{\partial v}{\partial t}=\frac{\delta E}{\delta v}=\nabla\cdot\Bigl(\varphi^{\prime}(v)\nabla v\Bigr)+\mu\left(t\right)v.$
Here, the variational derivation recovers the exact form of the flux term derived from conservation laws. We observe that the diffusive term $\Delta\varphi(v)$ can be expressed in the form:
$\Delta\varphi(v)=\nabla\cdot\Bigl(\varphi^{\prime}(v)\nabla v\Bigr).$
By choosing the specific function $\varphi(v)=v^{\alpha}$, the equation (1.1) can be obtained after rearrangement. The interpretations are as follows: the diffusive term $\Delta v^{\alpha}$ describes the dispersal of individuals in a habitat (representing the way in which the population spreads out in space), whereas the term $\mu\left(t\right)v$ represents the effect of population growth or mortality, modeling the birth and death processes. The boundary conditions (1.2) and the initial condition (1.3) complement the model by ensuring that the initial population distribution is predefined and that there is no flux of individuals across the boundary of the domain.
5. Cross-Disciplinary Relevance
The Bernoulli localization principle developed in this paper is not confined to a single application domain. In this section we show how Equation (1.1) and its structural decomposition via the Bernoulli ODE are relevant to physics, biology, economics, and engineering. For each discipline we identify the physical model, specify what the variables represent, explain how the localization principle applies, and state a concrete, verifiable prediction.
5.1. Physics: thermal diffusion in porous media
In the physics of porous media, the temperature $v(x,t)$ in a material with temperature-dependent thermal conductivity $k(v)=\alpha v^{\alpha-1}$ satisfies the nonlinear heat equation
$\frac{\partial v}{\partial t}=\nabla\cdot(k(v)\nabla v)+Q(t)\,v=\Delta v^{\alpha}+Q(t)\,v,$ (5.1)
where $Q(t)$ is a time-dependent volumetric heat source (or sink). This is precisely Equation (1.1) with $\mu(t)=Q(t)$. The power-law dependence $k(v)=\alpha v^{\alpha-1}$ arises in radiative heat transfer ($\alpha=4$, Stefan–Boltzmann regime) and in the thermal conductivity of certain polymers and ceramics [2, 12].
Verifiable prediction. For a porous medium with thermal conductivity $k(v)=\alpha v^{\alpha-1}$ (e.g., $\alpha=4$ for radiation-dominated transfer), initial temperature profile $v_{0}(x)$, and time-dependent heat source $Q(t)=Q_{0}/(1+t)$, the Bernoulli localization principle predicts that the temperature at any point $x$ and time $t$ satisfies
$S(t)\,u_{-}^{1/\alpha}(x)\leq v(x,t)\leq S(t)\,u_{+}^{1/\alpha}(x),$
where $u_{-}$ and $u_{+}$ are the elliptic sub- and supersolutions on the spatial domain, and $S(t)$ is computed from the closed-form expression (1.4) with $\mu(t)=Q_{0}/(1+t)$. These bounds can be tested against numerical simulations or experimental measurements of heat flow in porous ceramics.
5.2. Biology: population dynamics and tumor growth
Equation (1.1) was originally introduced by Gurtin and MacCamy [16] to model the dispersal of biological populations with density-dependent diffusion. In this context, $v(x,t)$ represents the population density, $\alpha>1$ captures the fact that organisms in crowded regions disperse more rapidly (the “slow diffusion” regime, where the diffusion coefficient vanishes when the density is zero, leading to compact support and finite-speed propagation), and $\mu(t)$ represents a time-varying net reproduction rate.
A particularly important biological application is tumor growth modeling. The density of tumor cells in a tissue region $\Omega$ evolves according to a porous-medium-type equation, where the power-law diffusion $\Delta v^{\alpha}$ captures the mechanical pressure exerted by the growing tumor mass, and $\mu(t)$ represents a time-varying proliferation rate that may decrease due to nutrient depletion or therapeutic intervention [18].
Verifiable prediction. For a tumor with initial density profile $v_{0}(x)$ and proliferation rate $\mu(t)=\mu_{0}e^{-\beta t}$ (representing exponential decay due to chemotherapy), the maximum tumor cell density at time $t$ is bounded above by $S(t)\max_{x}u_{+}^{1/\alpha}(x)$, where $S(t)$ is the explicit Bernoulli solution (1.4) with the given $\mu(t)$. This prediction can be tested against in vitro spheroid growth data.
5.3. Economics and social sciences: diffusion of innovations
The spread of a new technology, product, or idea through a population can be modeled by nonlinear diffusion equations [23]. In this framework, $v(x,t)$ represents the density of adopters at location $x$ and time $t$, the power-law diffusion $\Delta v^{\alpha}$ captures the nonlinear “peer-influence” effect (adoption accelerates when adopter density is high), and $\mu(t)$ represents a time-varying exogenous stimulus (e.g., advertising, policy incentives, or media coverage).
The domain $\Omega$ represents a geographical region, and the Dirichlet boundary condition $v=0$ on $\partial\Omega$ models the fact that adoption does not penetrate beyond the market boundary. The Bernoulli localization principle provides explicit bounds on the adoption density at any time and location.
Verifiable prediction. For a technology diffusion process with power-law peer influence ($\alpha=2$) and time-varying advertising intensity $\mu(t)=\mu_{0}/(1+t)$ (diminishing returns), the total adoption level $\int_{\Omega}v(x,t)\,dx$ is bounded between $S(t)\int_{\Omega}u_{-}^{1/2}(x)\,dx$ and $S(t)\int_{\Omega}u_{+}^{1/2}(x)\,dx$, where $S(t)$ is given by (1.4). These bounds can be tested against empirical adoption curves for products such as smartphones, electric vehicles, or renewable energy technologies.
5.4. Engineering: groundwater flow and materials processing
In hydrology, the hydraulic head $v(x,t)$ in an unconfined aquifer with nonlinear permeability satisfies a porous medium equation of the form (1.1), where $\alpha$ depends on the soil type and $\mu(t)$ represents time-varying recharge or extraction rates [3]. The Bernoulli localization principle provides rigorous upper and lower bounds on the hydraulic head, which are useful for predicting well yields, assessing contamination risks, and designing groundwater management strategies.
In materials science, the sintering of powder compacts—a process by which loose particles consolidate into a dense solid under heat—involves density-dependent diffusion governed by equations of the form (1.1), where $v(x,t)$ represents the local density of the compact and $\mu(t)$ represents the time-dependent sintering temperature profile [12]. The localization bounds predict the achievable density range at any stage of the sintering process.
Verifiable prediction. For groundwater flow in a sandy aquifer ($\alpha\approx 2$) with seasonal recharge $\mu(t)=\mu_{0}(1+\cos(2\pi t))$, the hydraulic head at any monitoring well location $x$ is bounded by the explicit Bernoulli barriers. These bounds can be compared against field measurements from piezometric monitoring networks.
6. Image Denoising Application and Comparison with Perona–Malik
6.1. Motivation: from biological diffusion to image processing
From an applied perspective, one of the most striking consequences of the framework developed in this paper is its direct applicability to digital image processing. To our knowledge, this constitutes the first application of a porous-medium-type (Gurtin–MacCamy) diffusion equation to image denoising.
The connection between diffusion equations and image processing has been recognized since the seminal work of Perona and Malik [21, 22], who proposed anisotropic diffusion as a method for edge-preserving image smoothing. Their model, however, is based on a linear diffusion equation with a spatially varying, gradient-dependent diffusion coefficient. The Gurtin–MacCamy model, by contrast, employs a genuinely nonlinear power-law diffusion mechanism $\Delta v^{\alpha}$ that has fundamentally different mathematical properties: finite-speed propagation, degenerate parabolicity, and compact support. These properties make it naturally suited for image denoising, as the nonlinear diffusion automatically slows down near edges (where the intensity is low after appropriate preprocessing) and acts more aggressively in homogeneous regions.
6.2. Mathematical formulation for image denoising
We treat a grayscale image (or each color channel of a color image) as a function $v:\Omega_{h}\times[0,T]\rightarrow[0,1]$, where $\Omega_{h}\subset\mathbb{Z}^{2}$ is the discrete pixel grid. The denoising process is governed by the discrete version of the Gurtin–MacCamy equation:
$v_{i,j}^{n+1}=v_{i,j}^{n}+\Delta t\left[\Delta_{h}(v^{\alpha})_{i,j}-\mu(n\Delta t)\,v_{i,j}^{n}\right],$ (6.1)
where $\Delta_{h}$ denotes the standard five-point discrete Laplacian
$\Delta_{h}(u)_{i,j}=u_{i+1,j}+u_{i-1,j}+u_{i,j+1}+u_{i,j-1}-4u_{i,j}.$ (6.2)
In the denoising application, the term $\mu(t)\,v$ acts as a time-dependent damping that suppresses noise more aggressively at early stages (when the noise level is high) and less aggressively as the image converges to a clean state. The function $\mu(t)=0.3/(1+t)$ implements this strategy: strong initial absorption that decays over time.
6.3. Comparison with Perona–Malik anisotropic diffusion
The Perona–Malik (PM) model [21, 22] is the standard benchmark for edge-preserving image smoothing. It replaces the linear diffusion $\Delta v$ with the anisotropic operator
$\nabla\cdot\bigl(g(|\nabla v|)\,\nabla v\bigr),$ (6.3)
where $g(s)=\exp(-s^{2}/K^{2})$ is an edge-stopping function. The PM model reduces diffusion near edges (where $|\nabla v|$ is large) and maintains full diffusion in flat regions.
The key difference is that the GM model achieves edge preservation through a fundamentally different mechanism: the power-law nonlinearity $\Delta v^{\alpha}$ with $\alpha>1$ automatically reduces the diffusion coefficient when the intensity $v$ is small, without requiring explicit gradient estimation. This makes the GM model more robust to noise in the gradient computation, which is a well-known weakness of the PM approach at high noise levels.
6.4. Algorithmic description
The complete denoising pipeline is summarized in Algorithm 1.
Algorithm 1 Gurtin–MacCamy Image Denoising
0:  Clean image $I\in[0,1]^{M\times N\times 3}$, noise level $\sigma$, exponent $\alpha$, time step $\Delta t$, number of steps $K$, growth function $\mu(\cdot)$
0:  Denoised image $I_{\mathrm{den}}\in[0,1]^{M\times N\times 3}$
1:  Generate noisy image: $I_{\mathrm{noisy}}\leftarrow\mathrm{clip}(I+\sigma\cdot\mathcal{N}(0,1),\;0,\;1)$
2:  for each color channel $c\in\{R,G,B\}$ do
3:   $v\leftarrow I_{\mathrm{noisy}}[:,:,c]$
4:   for $n=0,1,\ldots,K-1$ do
5:    $v_{\alpha}\leftarrow v^{\alpha}$ (elementwise power)
6:    $L\leftarrow\Delta_{h}(v_{\alpha})$ (discrete Laplacian via (6.2))
7:    $v\leftarrow v+\Delta t\cdot(L-\mu(n\Delta t)\cdot v)$
8:    $v\leftarrow\mathrm{clip}(v,\;0,\;1)$
9:   end for
10:   $I_{\mathrm{den}}[:,:,c]\leftarrow v$
11:  end for
12:  Apply postprocessing (contrast stretch + gamma correction) to $I_{\mathrm{den}}$
13:  return $I_{\mathrm{den}}$
6.5. Python implementation: transparent explanation
We now present the complete Python implementation with a detailed explanation of each component. The goal is full transparency: every design choice is justified mathematically or algorithmically, so that the reader can reproduce, modify, and extend the code.
6.5.1 Imports and parameters
1import numpy as np
2import imageio.v2 as imageio
3import matplotlib.pyplot as plt
4from skimage.metrics import structural_similarity as ssim
5
6alpha_gm = 4
7dt_gm = 0.0118
8steps_gm = 400
9
10dt_pm = 0.2
11steps_pm = 40
12K_pm = 0.1
13
14gamma_correction = 1.05
15stretch_strength = 0.9
16brightness_boost = 1.0
17
18sigma = 0.18
19
20def mu(t):
21 return 0.3 / (1 + t)
Listing 1: Imports and model parameters.
Explanation. The code imports four libraries: numpy for array operations, imageio for reading image files, matplotlib for visualization, and skimage.metrics for the SSIM quality metric.
The Gurtin–MacCamy parameters are: $\alpha=4$ (the diffusion exponent, corresponding to the radiation-dominated regime), $\Delta t=0.0118$ (the Euler time step, chosen small enough for stability of the explicit scheme), and $K=400$ (the number of diffusion steps). The Perona–Malik parameters are: $\Delta t=0.2$, $K=40$, and $K_{\mathrm{pm}}=0.1$ (the edge-stopping threshold). The postprocessing parameters control contrast recovery after denoising (see below).
The noise level $\sigma=0.18$ specifies the standard deviation of the additive Gaussian noise. The function $\mu(t)=0.3/(1+t)$ implements the time-dependent growth rate from the theoretical framework: it provides strong damping at early times (when noise is dominant) and progressively weakens as the image approaches a clean state.
6.5.2 Image quality metrics
1def compute_mse(a, b):
2 return np.mean((a - b)**2)
3
4def compute_psnr(a, b):
5 mse = compute_mse(a, b)
6 if mse == 0:
7 return float("inf")
8 return 20 * np.log10(1.0 / np.sqrt(mse))
9
10def compute_ssim_color(a, b):
11 return np.mean([
12 ssim(a[:,:,c], b[:,:,c], data_range=1.0)
13 for c in range(3)
14 ])
Listing 2: Quality metrics: MSE, PSNR, and SSIM.
Explanation. Three standard image quality metrics are implemented.
  • Mean Squared Error (MSE): $\mathrm{MSE}=\frac{1}{MN}\sum_{i,j}(a_{i,j}-b_{i,j})^{2}$, measuring the average squared pixel difference. Lower values indicate better reconstruction.
  • Peak Signal-to-Noise Ratio (PSNR): $\mathrm{PSNR}=20\log_{10}(1/\sqrt{\mathrm{MSE}})$ dB, measuring the ratio of peak signal to noise on a logarithmic scale. Higher values indicate better quality; a difference of 3 dB corresponds to a factor-of-two improvement in MSE.
  • Structural Similarity Index (SSIM): A perceptual metric that compares luminance, contrast, and structural information between two images. Values range from 1-1 to $1$, with $1$ indicating perfect similarity. For color images, the SSIM is computed per channel and averaged.
6.5.3 Postprocessing
1def postprocess(u):
2 v = u.copy()
3 v_min = v.min()
4 v_max = v.max()
5 if v_max > v_min:
6 v = (v - v_min) / (v_max - v_min)
7 v = stretch_strength * v + (1 - stretch_strength) * u
8 v = np.power(v, gamma_correction)
9 v *= brightness_boost
10 return np.clip(v, 0, 1)
Listing 3: Postprocessing: contrast stretching and gamma correction.
Explanation. After the diffusion process, the image intensities may have shifted and compressed. The postprocessing restores visual quality through three operations:
  1. 1.
    Contrast stretching: The image is linearly mapped to the full $[0,1]$ range by $v\mapsto(v-v_{\min})/(v_{\max}-v_{\min})$. This compensates for the intensity reduction caused by the damping term $\mu(t)\,v$ in the diffusion equation.
  2. 2.
    Blending: The stretched image is blended with the original diffused image using the parameter stretch_strength=0.9=0.9:
    $v_{\mathrm{out}}=0.9\cdot v_{\mathrm{stretched}}+0.1\cdot v_{\mathrm{diffused}}.$
    This prevents over-enhancement of low-contrast regions.
  3. 3.
    Gamma correction: A mild gamma correction ($\gamma=1.05$) is applied to adjust the luminance curve, slightly darkening the midtones for a more natural appearance.
6.5.4 The Gurtin–MacCamy diffusion kernel
1def diffuse_channel_gm(u, alpha, dt, steps):
2 u = u.copy()
3 for n in range(steps):
4 u_alpha = u**alpha
5 lap = (
6 np.roll(u_alpha, 1, axis=0) +
7 np.roll(u_alpha, -1, axis=0) +
8 np.roll(u_alpha, 1, axis=1) +
9 np.roll(u_alpha, -1, axis=1) -
10 4*u_alpha
11 )
12 u = u + dt * (lap - mu(n*dt)*u)
13 u = np.clip(u, 0, 1)
14 return u
Listing 4: Gurtin–MacCamy nonlinear diffusion for a single color channel.
Explanation. This function implements the core of the Gurtin–MacCamy denoising algorithm, corresponding to the discrete evolution equation (6.1). At each time step $n$:
  1. 1.
    Compute $v^{\alpha}$: The elementwise power $u^{\alpha}$ is computed. For $\alpha=4$, this strongly attenuates low intensities (noise) while preserving high intensities (signal).
  2. 2.
    Compute the discrete Laplacian: The five-point stencil (6.2) is implemented via np.roll, which shifts the array along each axis. The sum of four shifted copies minus $4$ times the original gives the standard finite-difference approximation to $\Delta(v^{\alpha})$. The use of np.roll implements periodic boundary conditions, which is standard practice in image processing and avoids boundary artifacts.
  3. 3.
    Euler update: The forward Euler step
    $v^{n+1}=v^{n}+\Delta t\bigl[\Delta_{h}(v^{\alpha})-\mu(n\Delta t)\,v^{n}\bigr]$
    advances the solution by one time step. The damping term $\mu(n\Delta t)\,v^{n}$ implements the time-dependent absorption from the Gurtin–MacCamy framework.
  4. 4.
    Clip to $[0,1]$: Pixel values are clamped to the valid range to maintain physical meaningfulness and numerical stability.
The mathematical connection to the theoretical framework is direct: this function implements a forward Euler discretization of the PDE $\partial_{t}v=\Delta v^{\alpha}-\mu(t)v$, which corresponds to equation (1.1) with the sign of the reaction term adapted for the absorption regime appropriate to denoising.
6.5.5 The Perona–Malik diffusion kernel
1def perona_malik_channel(u, dt, steps, K):
2 u = u.copy()
3 for _ in range(steps):
4 ux_f = np.roll(u, -1, axis=1) - u
5 ux_b = u - np.roll(u, 1, axis=1)
6 uy_f = np.roll(u, -1, axis=0) - u
7 uy_b = u - np.roll(u, 1, axis=0)
8
9 cxf = np.exp(-(ux_f**2) / (K**2))
10 cxb = np.exp(-(ux_b**2) / (K**2))
11 cyf = np.exp(-(uy_f**2) / (K**2))
12 cyb = np.exp(-(uy_b**2) / (K**2))
13
14 div = (cxf*ux_f - cxb*ux_b + cyf*uy_f - cyb*uy_b)
15 u = u + dt * div
16 u = np.clip(u, 0, 1)
17 return u
Listing 5: Perona–Malik anisotropic diffusion for a single color channel.
Explanation. This function implements the classical Perona–Malik anisotropic diffusion model [21, 22] for comparison. At each time step:
  1. 1.
    Compute finite differences: Forward and backward differences are computed along both spatial axes (horizontal: $x$, vertical: $y$), giving four gradient estimates per pixel.
  2. 2.
    Edge-stopping function: The Perona–Malik conductivity $g(s)=\exp(-s^{2}/K^{2})$ is applied to each gradient estimate. When the gradient is large (near an edge), $g\approx 0$ and diffusion is suppressed; when the gradient is small (in a flat region), $g\approx 1$ and full diffusion occurs.
  3. 3.
    Divergence and update: The discrete divergence of the flux $g\cdot\nabla v$ is computed and the Euler step is applied, as in the continuous equation $\partial_{t}v=\nabla\cdot(g(|\nabla v|)\nabla v)$.
The Perona–Malik model requires choosing the threshold $K$ carefully: too small, and noise is preserved; too large, and edges are blurred. This sensitivity is a recognized limitation of the PM approach, particularly at high noise levels where gradient estimates are unreliable.
6.5.6 The complete denoising pipeline
1img = imageio.imread("lenna.png").astype(np.float32) / 255.0
2gauss = sigma * np.random.randn(*img.shape)
3noisy_img = np.clip(img + gauss, 0, 1)
4
5den_gm = np.zeros_like(img)
6den_pm = np.zeros_like(img)
7
8for c in range(3):
9 den_gm[:,:,c] = diffuse_channel_gm(
10 noisy_img[:,:,c], alpha_gm, dt_gm, steps_gm
11 )
12 den_pm[:,:,c] = perona_malik_channel(
13 noisy_img[:,:,c], dt_pm, steps_pm, K_pm
14 )
15
16den_gm_pp = postprocess(den_gm)
17den_pm_pp = postprocess(den_pm)
18
19for label, im in [("Noisy", noisy_img),
20 ("Gurtin-MacCamy", den_gm_pp),
21 ("Perona-Malik", den_pm_pp)]:
22 print(f"{label}: MSE={compute_mse(img, im):.6f}, "
23 f"PSNR={compute_psnr(img, im):.2f} dB, "
24 f"SSIM={compute_ssim_color(img, im):.4f}")
Listing 6: Main pipeline: noise addition, denoising, and metric computation.
Explanation. The main pipeline operates in four stages:
  1. 1.
    Image loading and normalization: The test image (lenna.png) is loaded and normalized to $[0,1]$ by dividing by $255$.
  2. 2.
    Noise addition: Additive white Gaussian noise with standard deviation $\sigma$ is generated and added to the clean image. The result is clipped to $[0,1]$ to maintain valid pixel values.
  3. 3.
    Channel-wise denoising: Both the GM and PM denoising algorithms are applied independently to each of the three color channels (R, G, B). This channel-independent approach is standard in color image processing and avoids introducing color artifacts.
  4. 4.
    Postprocessing and evaluation: The postprocessing function restores contrast to the denoised images, and the three quality metrics (MSE, PSNR, SSIM) are computed for each method by comparing against the original clean image.
6.6. Results and comparison
The performance of our model, along with the metrics for the noisy image and the Perona–Malik results, is summarized in Table 1 for two different noise levels ($\sigma=0.1$ and $\sigma=0.18$).
Table 1: Comparative performance metrics for image denoising on lenna.png.
Noise Level Metric Noisy Image Gurtin–MacCamy Perona–Malik
$\sigma=0.1$ MSE 0.009564 0.002459 0.002529
PSNR (dB) 20.19 26.09 25.97
SSIM 0.2960 0.6583 0.6355
$\sigma=0.18$ MSE 0.028305 0.004723 0.015066
PSNR (dB) 15.48 23.26 18.22
SSIM 0.1533 0.5249 0.2226
6.7. Discussion of the denoising results
The results in Table 1 reveal several important observations.
Low noise ($\sigma=0.1$). At moderate noise levels, both the Gurtin–MacCamy (GM) and Perona–Malik (PM) models achieve substantial denoising, with the GM model holding a slight advantage across all three metrics: MSE improves from $0.009564$ (noisy) to $0.002459$ (GM) versus $0.002529$ (PM); PSNR increases from $20.19$ dB to $26.09$ dB (GM) versus $25.97$ dB (PM); and SSIM increases from $0.2960$ to $0.6583$ (GM) versus $0.6355$ (PM). The advantage of the GM model, while modest, is consistent.
High noise ($\sigma=0.18$). At elevated noise levels, the superiority of the GM model becomes dramatic. The GM model achieves PSNR =23.26=23.26 dB and SSIM =0.5249=0.5249, while the PM model achieves only PSNR =18.22=18.22 dB and SSIM =0.2226=0.2226—a difference of over $5$ dB in PSNR and more than double the SSIM. The PM model’s performance degrades sharply because its gradient-based edge-stopping function $g(|\nabla v|)$ becomes unreliable when the noise level is comparable to the gradient magnitudes of true edges. The GM model, by contrast, does not rely on gradient estimation; its edge-preserving capability arises from the intrinsic nonlinearity of the power-law diffusion, which is robust to additive noise.
The role of the time-dependent growth rate. The function $\mu(t)=0.3/(1+t)$ plays a crucial role in the GM model’s success. At early times, $\mu(t)$ is relatively large, providing strong absorption that suppresses noise aggressively. As time progresses, $\mu(t)$ decreases, allowing the diffusion to refine structural details without over-smoothing. This adaptive behavior, which emerges naturally from the theoretical framework of Equation (1.1), is a distinctive advantage of the GM approach.
The Python code implementing this comparison is publicly available at [11] for full reproducibility.
7. Conclusions and Outlook
7.1. Summary of contributions
In this paper, we have introduced the Bernoulli localization principle, a genuinely new approach to the analysis of nonlinear diffusion equations of porous medium type with time-dependent growth rates. The principle provides a complete structural description of solutions by explicitly decoupling temporal dynamics—governed by a Bernoulli-type ODE with a closed-form solution—from spatial profiles determined by a stationary sublinear elliptic equation. This decomposition changes the way the mathematical community can view this class of partial differential equations: rather than treating the parabolic PDE as an opaque object accessible only through qualitative comparison principles and energy estimates, the localization principle reveals its internal structure in explicit, computable terms.
The main theoretical contributions are:
  1. 1.
    Theorem 1.1: We proved the existence and uniqueness of separable classical solutions of the form
    $v(x,t)=S(t)u^{1/\alpha}(x),$
    where the spatial profile $u(x)$ satisfies a sublinear elliptic equation and the temporal function $S(t)$ is governed by a Bernoulli-type ODE. The separation-of-variables approach reduces the infinite-dimensional parabolic PDE to a scalar ODE plus a stationary elliptic problem, yielding explicit analytical expressions that are directly implementable.
  2. 2.
    Theorem 1.2: We established a comparison principle showing that if the initial population $v_{0}$ is smaller than $w_{0}$ and the growth rate $\mu_{1}(t)$ is smaller than $\mu_{2}(t)$, then the solution $v(x,t)$ remains bounded by $w(x,t)$ for all subsequent times.
  3. 3.
    Theorem 1.3: We extended the existence and uniqueness theory to the general non-separable case using the monotone iteration method based on subsolution–supersolution techniques, where the sub- and supersolutions are explicitly constructed from the Bernoulli ODE solution. This provides concrete, quantitatively sharp localization bounds for every solution whose initial data lie in the ordered interval $v_{-},v_{+}]$.
  4. 4.
    Image denoising: We presented the first application of a porous-medium-type diffusion equation to digital image processing, showing that the Gurtin–MacCamy model with time-dependent growth rate achieves denoising performance comparable to or exceeding the classical Perona–Malik filter, particularly at high noise levels.
7.2. Impact and broader significance
The Bernoulli localization principle changes the paradigm for analyzing porous-medium-type diffusion equations in several ways:
  • It provides a structural decomposition that was not previously available, making explicit what was formerly implicit.
  • It converts qualitative comparison principles into quantitative localization bounds with explicitly computable barriers.
  • It reduces the computational complexity of solving the parabolic PDE by decoupling the temporal and spatial components.
  • It enables direct parameter sensitivity analysis, since the dependence on $\alpha$, $\gamma$, and $\mu(\cdot)$ is transparently readable from the closed-form expressions.
The cross-disciplinary relevance of the framework has been demonstrated through concrete models and verifiable predictions in physics (thermal diffusion in porous media), biology (population dynamics, tumor growth), economics (diffusion of innovations), and engineering (groundwater flow, materials processing). In each case, the Bernoulli localization principle provides quantitative predictions that can be tested against data.
7.3. Future research directions
Several directions for future work are natural:
  1. 1.
    Higher-dimensional domains and complex geometries: Extension of the numerical implementation to three-dimensional domains with irregular boundaries, using finite element discretizations of the elliptic problem.
  2. 2.
    Spatially varying diffusion coefficients: Generalization of the localization principle to equations of the form $\partial_{t}v=\nabla\cdot(a(x)\nabla v^{\alpha})+\mu(t)v$, where $a(x)$ models spatial heterogeneity.
  3. 3.
    Nonlinear boundary conditions: Extension to problems with nonlinear flux conditions on the boundary, which arise in models of biological membranes and thermal radiation.
  4. 4.
    Stochastic perturbations: Analysis of the robustness of the localization principle under stochastic forcing, relevant to models with environmental noise.
  5. 5.
    Advanced image processing: Exploration of the GM diffusion model for other image processing tasks such as segmentation, inpainting, and super-resolution, potentially in combination with machine learning techniques.
  6. 6.
    Experimental validation: Comparison of the theoretical predictions with experimental data from laboratory measurements of diffusion in porous materials, biological growth assays, or groundwater monitoring networks.
Acknowledgments
The author thanks the anonymous referees for their careful reading and helpful suggestions that improved the presentation of this paper.
Funding
This research received no external funding.
Data Availability
The Python code for all numerical experiments is publicly available at the URLs listed in the references. The test image used for the denoising experiments (lenna.png) is a standard benchmark image widely available in the image processing community.
Declarations
Conflicts of Interests: The author declares no conflict of interest.